### v1

# Actor-critic algorithms for fiber sampling problems

Abstract
We propose an actor-critic algorithm for a family of complex problems arising in algebraic statistics and discrete optimization. The core task is to produce a sample from a finite subset of the non-negative integer lattice defined by a high-dimensional polytope. We translate the problem into a Markov decision process and devise an actor-critic reinforcement learning (RL) algorithm to learn a set of good moves that can be used for sampling. We prove that the actor-critic algorithm converges to an approximately optimal sampling policy.
To tackle complexity issues that typically arise in these sampling problems, and to allow the RL to function at scale, our solution strategy takes three steps: decomposing the starting point of the sample, using RL on each induced subproblem, and reconstructing to obtain a sample in the original polytope. In this setup, the proof of convergence applies to each subproblem in the decomposition.
We test the method in two regimes. In statistical applications, a high-dimensional polytope arises as the support set for the reference distribution in a model/data fit test for a broad family of statistical models for categorical data. We demonstrate how RL can be used for model fit testing problems for data sets for which traditional MCMC samplers converge too slowly due to problem size and sparsity structure. To test the robustness of the algorithm and explore its generalization properties, we apply it to synthetically generated data of various sizes and sparsity levels.
1 Introduction
A variety of problems in computational mathematics and statistics are supported on discrete sets whose sizes explode combinatorially with the dimension of the problem. A typical example arises in discrete optimization problems, for which feasible solutions are integer points in a polytope. The polyhedral system $\mathcal{P}_{M}(b):=\{x\in\mathbb{R}^{d}:Mx=b,x\geq 0\}$, with $M\in\mathbb{Z}^{n\times d}$, is a standard form representation of any system of polynomial equations and inequalities. We are interested in the general problem of sampling integer points in $\mathcal{P}_{M}(b)$. Following the algebraic statistics literature [17], we will refer to the set $\mathbb{Z}^{d}\cap\mathcal{P}_{M}(b)$ as the $b$-fiber of $M$. Due to its universality, it is well understood that exploration of lattice points in a polytope is an NP-hard problem. A prominent application is testing statistical model fit to discrete data. In statistics, performing the test of model fit to the data is a fundamental task, necessary for the correct use of statistical models in a broad set of applications such as cross-classified census data, social networks, or neuronal networks. Yet for most categorical datasets in practice, the sheer size of the problem–or sparsity in the data–render all available computational tools practically infeasible. The challenge lies in computationally determining the tests’ reference distribution, which is supported on a prohibitively large and high-dimensional fiber.
A brief look back into the history of sampling discrete structures suggests that randomized search algorithms could simultaneously tackle sparsity, dimensionality, and time complexity of the problem, given their long history of successfully solving NP-hard problems. The Monte Carlo method [35] has found an application in almost every problem involving sampling. Its combination with the theory of Markov chains, the Markov chain Monte Carlo (MCMC) method has lead to an explosion of randomized search algorithms such as [34, 22, 18]. Although MCMC has been around for decades, algorithm design for fiber exploration is relatively young. Standard statistics theory offers asymptotic solutions to the testing problem, yet they often do not apply in practice. This has been known since the late 1980s [21]; the fundamental issue necessitates a synthetic data sample from a fiber associated to the statistical model. Theoretical guarantees for constructing irreducible Markov chains on fibers were provided in the seminal paper [15]; while Besag and Clifford argue in contemporaneous work [8] that it is best to produce an exchangeable sample from the fiber and not rely on MCMC, which may be slow to converge and produce a good enough sample for testing. The last three decades have seen an explosion of research in this area inspired by [15], see [2] for a summary of the state of the art. This brings us to our starting point: even with randomization, in applications dealing with high-dimensional polytopes $\mathcal{P}_{M}(b)$, there is a computational bottleneck for problems at scale. Unfortunately, this can already happen when the design matrix $M$ has just hundreds of columns.
Another line of research that has taken off since the 1950s is reinforcement learning (RL) and optimal control, pioneered by Richard Bellman [6]. Reinforcement learning algorithms can be thought of as informed random walks, in which the agent–or controller–learns from previous experiences through a reward system, in the process refining her understanding of the underlying problem dynamics. Every RL algorithm operates on a Markov decision process (MDP) or partially observed MDP. In their simplest form, MDPs are characterized by a set of states, actions, and corresponding rewards. The goal of an agent is to optimize a given objective, also known as the state-value function of the MDP. In this process, the agent learns the optimal policy, or control, which is a function mapping states to actions. In general, RL algorithms can be split into actor (policy-based), critic (value-based), and actor-critic (policy-value-based); further, they can be classified as on-policy or off-policy algorithms. Some of the earliest critic algorithms include policy and value iterations [25] which use iterative updates to refine the agent’s policy and state-value function. Another common family of value-based algorithms uses temporal differences (TD), computed by the Bellman equation. A representative algorithm of this family is Q-learning developed by Watkins [48], which uses an action-value instead of a state-value function; other notable work includes [43] and [47]. All of these try to learn a certain score function first, and then subsequently extract the optimal policy. In contrast, in actor algorithms, the agent seeks to directly find the best policy which optimizes the given objective. Policies are usually parameterized by neural networks and the parameters are updated through gradient descent [45]. Other parametrizations can be found in [19], [26], [33], and [49]. What makes such algorithms attractive is that they do not suffer from the curse of dimensionality, as neural networks can handle high dimensional data. Finally, actor-critic algorithms [5] improve performance by taking advantage of both appoaches: they use both the parametrization of the policy and temporal difference updates of the state-value function to solve the given MDP. Notable work in this area includes [32], [47].
The main theoretical contribution of this work is a general solution to a family of difficult fiber sampling problems with theoretical guarantees. Specifically:
  1. 1.
    This manuscript designs an actor-critic fiber sampling scheme, building on the existing A2C algorithm proposed by [36]. Our modified version of the algorithm uses the same feature extractor for both actor and critic networks, hence reducing the number of parameters.
  2. 2.
    We provide the proof of asymptotic convergence to the optimal policy. To the best of our knowledge, this is the first proof of convergence for an actor-critic algorithm using (1) a generalized advantage estimator, first introduced in [40], and (2) the shared feature extractor.
  3. 3.
    In order for the agent to be able to move through a very high-dimensional and potentially sparse fiber, we use linear–rather than non-linear–algebraic computation as input to the action space determination. We also discuss an additional step to decompose the initial solution and allow the agent to learn on sub-problems. This decomposition step further alleviates the computational burden of producing moves necessary for fiber sampling.
To illustrate the performance of our method on real data, we use a statistical application. This setting is the natural sandbox, in which we can address the practical challenges that have remained open for sampling certain fibers at scale. We showcase how the algorithm can produce an exchangeable sample of synthetic networks in a high-dimensional sparse fiber for a data set of applied interest in network analysis [27]. While thoroughly demonstrating correctness of statistical conclusions for various data sets is out of scope for this paper (after all, testing large data sets is an open problem!), we note that statistical conclusions arising from our RL-generated sample do match those found in [29]. Finally, we generate synthetic data of various sizes and sparsity levels, demonstrate algorithm performance, and discuss further work on optimizing the decomposition step.
2 Related literature
Machine learning has demonstrated success in solving or improving heuristics for combinatorial optimization problems. For example, [46] presents an RL approach to the cutting plane method used in all modern integer programming solvers. A thorough overview and perspective of applicability and impact of RL in combinatorial optimization is provided in [7]. Very recently, [51] proposed casting a famous NP-complete problem into a Markov Decision Process and applying a specific reinforcement learning algorithm to solve it: they focus on the general integer feasibility problem, whether a linear system of equations and inequalities has an integer solution. It is instructive to draw parallels and highlight differences in our work compared to [51], which serves as our the starting point in terms of algorithm design. In the interest of space, we offer a detailed discussion on this in Appendix A.1.
To understand the impact RL can have on fiber sampling in practice, it is helpful to also reflect on some of the literature for fiber sampling from the point of view of statistical applications. The biggest drawback of the Metropolis-Hastings algorithm proposed for goodness-of-fit testing in [15] is that it requires one to compute a Markov basis [37] of the design matrix $M$ defining the polytope $P_{M}(b)$. However, computing the full Markov basis for large problems is simply impossible. Building on recent progress in the statistics literature [28, 20], [4] uses the smaller lattice basis as a starting set of moves and a biased sampling scheme in its random exploration in order to guide the search towards denser regions of the fiber; with careful parameter tuning, it could also sample relatively sparse fibers. There is a broader literature on fiber sampling based on, e.g., sequential importance sampling (SIS). Note that [16] argues that the Markov bases approach computed dynamically performed better than SIS (see also [4, Remark 2.2]). In the present paper, we focus on using RL to address the shortcomings and scaling issues for sampling approaches based on Markov bases and sub-bases.
3 Fiber Sampling as a Markov Decision Process (MDP)
Fiber basics. Fix the integer design matrix $M\in\mathbb{Z}^{n\times d}$, and one lattice point $u\in\mathbb{Z}_{\geq 0}^{d}$, and consider the set of integer points in the polytope $\mathcal{P}_{M}(b)$ such that $b=Mu$. We will use the following notation for the fiber of $M$: $\mathcal{F}_{M}(u):=\{v\in\mathbb{Z}_{\geq 0}^{d}\leavevmode\nobreak\ |Mu=Mv\}.$ Rather than fixing $b$ and then searching for a feasible point $u$ for the system, we start with a given data point $u$ and consider the fiber it defines. This mimics the situation in statistics, where $u$ is the observed categorical data and $M$ defines the statistical log-linear model [17]. The vector $Mu$ is the observed value of the sufficient statistics for the model corresponding to the observed data $u$; oftentimes in applications, $u$ will represent a vector flattening of a contingency table, which is a $k$-dimensional tensor of format $d_{1}\times\dots,\times d_{k}$ with integer entries, in which case $d:=d_{1}\cdots d_{k}$. This is the setting for the application in Section 6.
What it takes to move. Any two points $v_{1},v_{2}\in\mathcal{F}_{M}(u)$ satisfy $M(v_{1}-v_{2})=0$. Any vector $v_{1}-v_{2}\in\ker M$ is called a move on the fiber. The vector space basis of the lattice $\ker M$ is any set of $n-rank(M)$ vectors in $\ker M$ that are linearly independent; we call any such spanning set the lattice basis of $M$. For a summary of the many bases of the integer lattice $\ker M$, the reader is invited to consult [17, §1.3]. We view any lattice basis as a set of moves that connects the fiber as follows:
Definition 3.1 (Lattice basis).
A lattice basis of a design matrix $M$ is any set of $c:=n-rank(M)$ linearly independent moves $\mathcal{B}(M):=\{l_{1},\dots,l_{c}\}\subset\mathbb{Z}^{d}$, $Ml_{i}=0$. For any data $u\in\mathbb{Z}_{\geq 0}^{d}$ and for any $v\in\mathcal{F}_{M}(u)$, there exist $l_{i_{1}},\dots,l_{i_{k}}\in\mathcal{B}(M)$ and corresponding scalars $\alpha_{i_{1}},\dots,\alpha_{i_{k}}\in\mathbb{Z}$ such that the linear combination can be used to reach $v$ from $u$:  $u+\alpha_{i_{1}}l_{i_{1}}+\dots+\alpha_{i_{k}}l_{i_{k}}=v$.
Note that any integral linear combination of elements of $\mathcal{B}(M)$ is a move:
$\alpha_{i_{1}}l_{i_{1}}+\dots+\alpha_{i_{k}}l_{i_{k}}\in\ker M.$ (1)
Lattice bases alone cannot be used to directly create a random walk on the fiber due to the fact that any sub-sum in the above equation may result in negative entries in the vector. Enter Markov bases: any set of moves :=$\mathcal{B}:=\{b_{1},\dots,b_{m}\}\subset\mathbb{Z}^{d}$, $m\geq c$, $Mb_{i}=0$, such that for any data $u\in\mathbb{Z}_{\geq 0}^{d}$ and for any $v\in\mathcal{F}_{M}(u)$, there exist $b_{i_{1}},\dots,b_{i_{N}}\in\mathcal{B}$ that can reach $v$ from $u$ while being on the fiber:
$u+b_{i_{1}}+\dots+b_{i_{N}}=v,\mbox{ with the condition that }u+b_{i_{1}}+\dots+b_{i_{j}}\in\mathbb{Z}^{d}_{\geq 0}\mbox{ for all }j\leq N.$
The power of Markov bases is they exist and are finite for any integer matrix $M$ by the Hilbert basis theorem, and they guarantee a connected Markov chain on all fibers of $M$ simultaneously. This notion, introduced in [15], has generated close to a 1000 research papers to date. The following points are essential for us (cf. [37]): a Markov basis is the minimal set of moves that guarantees fiber connectivity, but it does so for every possible fiber of a given $M$, i.e., all possible observed data $u$. For a given $u$, one rarely needs all of the basis elements to connect that specific fiber [16]. Markov bases can be arbitrarily complicated, but there are bounds on their complexity in special cases; see [14], [2]. A lattice basis of $M$ is a minimal sub-basis in the sense of [13, Lemma 4.2]): $\mathcal{B}(M)$ generates the lattice as a vector space and the saturation of the corresponding ideal results in the full toric ideal; while the detailed algebraic discussions are out of scope for this work, the interested reader can find details in [41] and [15]. A Gröbner basis [42] of $M$ always contains a minimal Markov basis.
Recasting fiber moves into an RL framework. Any set of moves defined above can be used as a set of actions for exploring the fiber in the context of reinforcement learning. In terms of scalability, lattice bases are the only bases that can be computed for very large matrices $M$, because other sets require non-linear algebra computations. With this in mind, we define the MDP of interest.
Definition 3.2.
(Fiber Sampling Markov Decision Process) Fix an integer matrix $M\in\mathbb{Z}^{n\times d}$ and $s_{0}\in\mathbb{Z}_{\geq 0}^{d}$, a given initial solution determining the fiber $\mathcal{F}_{M}(b)$ with $b=Ms_{0}$. A Fiber Sampling Markov Decision Process with planing horizon $T=\infty$ is a stationary process on the set of data $(\mathcal{S},\mathcal{S}^{-},\mathcal{S}^{0},\mathcal{A},p,R_{t},\gamma)$, for $t=0,1,\dots$, defined as follows.
The state space $\mathcal{S}:=\mathcal{F}_{M}(b)$ consists of all integral points in the polytope $\mathcal{P}_{M}(b)$. The set of unfeasible states is $\mathcal{S}^{-}:=\{s\in\mathbb{Z}^{d}\ |\ Ms=b,s\not\geq 0\}$, while $\mathcal{S}^{0}$ is the set of discovered feasible states. At the start of the algorithm, $\mathcal{S}^{0}=\{s_{0}\}$.
$\mathcal{A}$ is the action space consisting of moves derived from the linear combination of lattice basis vectors, so that each action is given as $a=\alpha_{1}l_{1}+\dots\alpha_{c}l_{c}$, with $l_{i}\in\mathcal{B}(M)$. The random variable whose realization at time $t$ is action $a$ is denoted by $A_{t}$.
The transition kernel is deterministic; i.e., the transition dynamics are simply $S_{t+1}=S_{t}+A_{t}.$
The mapping $R_{t}:\ \mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the undiscounted reward at time $t$, and $\gamma$ is the discount factor. The reward function $R_{t+1}$, obtained after time $t$ is defined by two sub-rewards. Namely, $R_{t+1}(S_{t},A_{t},S_{t+1})=R^{feas}_{t+1}(S_{t},A_{t},S_{t+1})+R^{zero}_{t+1}(S_{t},A_{t},S_{t+1})$ where
$\displaystyle R^{feas}_{t+1}(S_{t},A_{t},S_{t+1})$ $\displaystyle=\sum_{i=1}^{d}I_{(S_{t}+A_{t})_{i}<0}(S_{t}+A_{t})_{i}$ (2)
$\displaystyle R^{zero}_{t+1}(S_{t},A_{t},S_{t+1})$ =$\displaystyle=\begin{cases}-d&\text{ if }A_{t}\equiv\textbf{0}\\ 0&\text{ otherwise }.\end{cases}$ (3)
The maximization objective is given by
$\max_{\pi}\mathbb{E}_{\pi}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}(S_{t},\pi(A_{t}|S_{t}),S_{t+1})\Big{]},$ (4)
where $\tau=(s_{0},a_{0},s_{1},a_{1},\dots)$ is a trajectory sampled under policy $\pi$. The probability of observing a trajectory $\tau\in\Omega$ is computed as:
$p(\tau)=p(S_{0})\pi(A_{0}|S_{0})\prod_{t=1}^{\infty}p(S_{t}|S_{t-1},A_{t-1})\pi(A_{t}|S_{t})=p(S_{0})\pi(A_{0}|S_{0})\prod_{t=1}^{\infty}\pi(A_{t}|S_{t}).$
We will denote a truncated finite trajectory by $\tau_{t:t+h}:=(s_{t},a_{t},\dots,s_{t+h-1},a_{t+h-1},s_{t+h})$.
In Equation (2), $I$ denotes the indicator function, so $R^{feas}_{t+1}$ penalizes the agent based on the magnitude of negative coordinates in $\mathcal{S}^{-}$, if the action vector is not feasible. On the other hand, $R^{zero}_{t+1}$ penalizes the agent for picking the zero action, which would not let the agent explore the environment.
Remark 3.1 (Action space in practice: bounded coefficients).
A critical problem [51] faced was the size of the action space. They used a lifting of the Gröbner basis of the 3-way transportation polytope specified by the linear embedding, and considered actions with entries in $\{-1,0,1\}$. Their agent’s policy $\pi$ produces the actions in the form of tables directly from a neural network.
In our case, we use linear combinations of lattice basis elements to produce actions. However, such set of moves is also very large, given the range of coefficients $\alpha_{i}$; the best universal upper bound on the $\alpha_{i}$ is too large (deriving, e.g., from [41, Theorem 4.7]). In this case, convergence of the RL algorithm will be extremely slow due to the fact that the agent has to try each action for each state she encounters. Therefore, we only consider moves with coefficients with pre-specified bounds:
$\forall i\in[N],\ \leavevmode\nobreak\ \min\alpha_{i}=c_{1}\in\mathbb{Z},\ \max\alpha_{i}=c_{2}\in\mathbb{Z},\text{ such that }c_{1}<c_{2}.$
Moreover, our agent’s policy only produces coefficients $\alpha_{i}$, which are then used to create a move as a linear combination of lattice bases as in Equation (1).
We design the Markov decision process in a way that when the agent finds herself in the unfeasible state $s\in\mathcal{S}^{-}$, all rewards onward are equal to zero. In turn, equation 4 becomes
$\displaystyle\max_{\pi}\mathbb{E}_{\pi}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}(S_{t},\pi(A_{t}|S_{t}),S_{t+1})\Big{]}=\max_{\pi}\mathbb{E}_{\pi}\Big{[}\sum_{t=0}^{T}\gamma^{t}R_{t+1}(S_{t},\pi(A_{t}|S_{t}),S_{t+1})\Big{]}+\sum_{t=T+1}^{\infty}0.$
It follows that the fiber sampling problem naturally breaks into an episodic task, each episode starting independently from a previously discovered state $s_{0}\in\mathcal{S}^{0}$, with assigned probability following a uniform distribution over all discovered states up to that time. Episode independence is important in de-correlating network training samples of states, actions and rewards across different episodes. Other technical details regarding the well-posedness of the objective, such as finiteness of the limit of the cumulative reward and ergodicity of the underlying policy and its stationary distribution, are presented in Appendices A.3 and A.2, respectively.
4 Actor-critic algorithm for fiber sampling
The policy. The policy $\pi$ is typically modeled by a parameterized function $\pi_{\theta}$. In our case, the parameterized policy is constructed out of two fully connected deep neural networks. The first part is called the feature extractor, it takes as input the current state $S_{t}=s$ at time $t$, and it tries to extract certain features shared by all elements in the state space. The second part is the policy network $h(s,a;\theta)$ with a soft-max function applied to it such that the output is given as the probability $\pi_{\theta}(a|s):=\frac{e^{h(a,s;\ \theta)}}{\sum_{s^{\prime}}e^{h(a,s^{\prime};\ \theta)}}$, over the action space. The part of the architecture in Figure 1 depicted by full lines refers to the policy, with $\phi_{\theta}$ representing the feature extractor.
Assumption 4.1.
The subset $\Theta\subset\mathbb{R}^{n}$ of network parameters is closed and bounded.
This is perhaps the strongest assumption we impose throughout this work. It is consistent with [24], which assumes compact support of networks over the parameter space. In practice, this is satisfied if one takes a large enough closed ball around initial parameters $\theta_{0}$.
Claim 4.1.
(a) For every $s\in\mathcal{S}$ and $a\in\mathcal{A}$,  $\pi_{\theta}(a|s)>0$. (b) For every $(s,a)\in\mathcal{S}\times\mathcal{A}$, the mapping $\theta\mapsto\pi_{\theta}(a|s)$ is twice continuously differentiable. Furthermore, the $\mathbb{R}^{n}$-valued function $\theta\mapsto\nabla_{\theta}\ln\pi_{\theta}(a|s)$ is bounded and has a bounded first derivative for any fixed $s$ and $a$.
That the gradient is bounded essentially follows from the policy definition above and finiteness of the state and action spaces. Detailed proof is in Appendix A.4.
The actor.
The minimization objective of the actor is given by the objective function $L_{a}(\theta)$, defined as $L_{a}(\theta):=V_{\pi_{\theta}}(s_{0})$, which is exactly the state-value function starting from the initial state $S_{0}=s_{0}$. Using the Policy Gradient Theorem [44], we can compute the gradient of the objective function as $\nabla_{\theta}L_{a}(\theta)=\nabla_{\theta}V_{\pi_{\theta}}(s_{0})=\mathbb{E}_{\pi_{\theta}}\Big{[}Q_{\pi_{\theta}}(S_{t},A_{t})\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}$. We can modify the gradient computation above in a way that the value of the gradient remains the same, while the variance of an estimator of this gradient is reduced. We accomplish this by introducing a baseline function, denoted $b(\cdot)$, which doesn’t depend on the action. The gradient is then $\nabla_{\theta}L_{a}(\theta)=\nabla_{\theta}V_{\pi_{\theta}}(s_{0})=\mathbb{E}_{\pi_{\theta}}\Big{[}(Q_{\pi_{\theta}}(S_{t},A_{t})-b(S_{t}))\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}$. A popular choice is $b(S_{t})=V_{\pi_{\theta}}(S_{t})$, which gives us the the definition of the advantage function $A_{\theta}(S_{t},A_{t}):=Q_{\pi_{\theta}}(S_{t},A_{t})-V_{\pi_{\theta}}(S_{t})$. This function measures how good or bad the action $A_{t}$ is at state $S_{t}$, compared to the policy $\pi_{\theta}$. We take the derivation one step further and introduce the Generalized Advantage Estimator $\widehat{A_{GAE}}(t)$ [40]; as the name suggests, it estimates the advantage function. Using $\widehat{A_{GAE}}(t)$ with Fubini-Tonelli theorem, in place of $A_{\theta}(S_{t},A_{t})$, we define the following policy gradient:
$\nabla_{\theta}L_{GAE}(\theta):=\mathbb{E}_{\pi_{\theta}}\Big{[}\widehat{A_{GAE}}(t)\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\mathbb{E}_{\pi_{\theta}}\Big{[}\delta(t+l)\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}.$
All the above computations have been derived under the assumption that we have an exact expression of the state-value function $V_{\pi_{\theta}}$. However, in practice, this is usually not the case and we can only work with an approximation of such a function under some parametrization $\omega$. In what follows, we show that such a parametrization does not pose any issue in the gradient estimate. We follow the work of [31] and [47] and define, for any $\theta$, an inner product $\langle\cdot,\cdot\rangle_{\theta}$ of two real-valued functions $F_{1},F_{2}$, on $\mathcal{S}$ and $\mathcal{S}\times\mathcal{A}$, viewed as vectors in $\mathbb{Z}^{|\mathcal{S}|}$ and $\mathbb{Z}^{|\mathcal{S}||\mathcal{A}|}$, by $\langle F_{1},F_{2}\rangle_{\theta}:=\sum_{s}\mu_{\theta}(s)\sum_{a}\pi_{\theta}(a|s)F_{1}(s)F_{2}(s,a)=\sum_{s}\mu_{\theta}(s)F_{1}(s)\sum_{a}\pi_{\theta}(a|s)F_{2}(s,a)$. Define a projection $\Pi_{\theta}:\mathbb{Z}^{|\mathcal{S}|}\mapsto\Psi_{\theta}$ as $\Pi_{\theta}F:=\arg\min_{\hat{F}\in\Psi_{\theta}}||F-\hat{F}||_{\theta}$. Here, $||\cdot||_{\theta}$ is the norm on $\mathbb{Z}^{|\mathcal{S}|}$ induced by the inner product $\langle\cdot,\cdot\rangle_{\theta}$. By definition of a projection, it follows that the error of the projection is orthogonal to the projection space. Hence, $\langle\Pi_{\theta}V_{\pi_{\theta}}-V_{\pi_{\theta}},\psi_{\theta}^{i}\rangle_{\theta}=0$ implies that $\langle\Pi_{\theta}V_{\pi_{\theta}},\psi_{\theta}^{i}\rangle_{\theta}=\langle V_{\pi_{\theta}},\psi_{\theta}^{i}\rangle_{\theta}$ which in turn shows that it suffices to learn the projection of $V_{\pi_{\theta}}$ onto $\Psi_{\theta}$ in order to compute the approximation of the gradient $g_{a}(\theta)$. Let $V_{\omega}$ be a projection of the state-value and let $\delta_{\omega}(t)$ denote the $t$-th temporal difference which uses the projected state-value function. Then we write the gradient approximation of the exact gradient as $g_{a}(\theta,\omega):=\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\mathbb{E}_{\pi_{\theta}}\Big{[}\delta_{\omega}(t+l)\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}$ which is estimated as $\hat{g_{a}}(\theta,\omega):=\frac{1}{K}\sum_{k=0}^{K-1}\nabla_{\theta}\ln\pi_{\theta}(A_{k}|S_{k})\sum_{l=k}^{K-1}(\gamma\lambda)^{l-k}\delta_{\omega}(t+l)$. Finally, the actor’s parameter update is given by the following.
Definition 4.1 (Actor update).
Let $\theta_{t}$ denote the policy parameter (actor) at time $t$. Then, we define the actor update along a truncated trajectory $\tau_{t:t+K}$ as
$\displaystyle\theta_{t+K}=\theta_{t+K-1}+\alpha(t+K-1)\Big{[}\frac{1}{K}\sum_{k=0}^{K-1}\nabla_{\theta}\ln\pi_{\theta}(A_{t+k}|S_{t+k})\sum_{l=k}^{K-1}(\gamma\lambda)^{l-k}\delta_{\omega}(t+l)\Big{]}.$ (5)
The update function above mathematically describes the update procedure of the stable-baseline3’s A2C actor code [38]; source code can be found at https://github.com/DLR-RM/stable-baselines3. For our algorithm, we slightly modify the library so that both policy and state value networks use shared feature extractor outputs, reflected in Figure 1.
The mathematical details of the derivation in this section can be found in Appendix A.5.
The critic. Not having access to the analytical form of state-value function $V$, we parametrize it:
$V_{\omega}(S_{t}):=\phi_{\theta}(S_{t})^{T}\omega\approx V_{\theta}.$
Assumption 4.2.
The subset $\Omega\subset\mathbb{R}^{m}$ of critic parameters is closed and bounded.
This assumption, similar to 4.1, is consistent with assumptions in [24]. In practice, this is satisfied if we take a large enough closed ball around initial parameters $\omega_{0}$.
The feature vector $\phi_{\theta}$ depends on the actor parameter $\theta$ and is shared between the actor and the critic. Such setup is a deviation from [47] and most of the other actor-critic algorithms. To the best of our knowledge, the convergence of algorithms that use this architecture has only been explored in [31]. Pictorially, the parameterization of the critic is given by the dashed pipeline in Figure 1, from the state $\mathcal{S}$, through the feature extractor parametrized by $\phi_{\theta}$, and finally to the critic network $\omega^{T}\phi_{\theta}$. The activation functions of the feature extractor are chosen such that the mapping $\theta\mapsto\phi_{\theta}$ is continuous, bounded and differentiable. The full claim with proof can be found in Appendix A.6.
To enable the critic to learn the approximate state-value function parametrized by $\omega$, we need to define a correct minimization objective for the critic. In our setup, the algorithm is learning the necessary weights in an on-line manner, meaning that the only observations that the algorithm can use to learn are the ones that are observed along some sampled trajectory. Work in [44] gives us several choices of minimization objectives which work with incremental data. One of the most basic but well established minimization objectives minimizes the difference between collected rewards along a truncated trajectory $\tau_{t:t+K}$ and the parametrized state-value function $V_{\omega}$. Focusing on the temporal difference in this objective, we replace it by a truncated version of the generalized advantage estimators; the detailed formulas appear in Appendix A.7. This leads to the following:
Definition 4.2 (Critic Minimization Objective).
The minimization objective for the critic is:
$\displaystyle L_{g_{c}}(\theta,\omega)=\min_{\omega}\Big{|}\Big{|}\frac{1}{2}\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega}(t+l)^{2}\Big{|}\Big{|}_{\theta}^{2},\ \text{ for }\omega\in\Omega,\theta\in\Theta.$ (6)
The update function above mathematically describes the update procedure of the stable-baseline3’s A2C critic code [38].
Definition 4.3 (Critic update).
Let $\omega_{t}$ denote the state-value parameter (critic) at time $t$. Then, we define the critic update along a truncated trajectory $\tau_{t:t+K}$ as
$\displaystyle\omega_{t+K}$ $\displaystyle=\omega_{t+K-1}+\beta(t+K-1)\Big{[}\frac{1}{K}\sum_{k=0}^{K-1}\Big{[}\sum_{l=k}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t+K-1}}(t+l)\nabla_{\omega_{t+K-1}}\delta_{\omega_{t+K-1}}(t+l)\Big{]}\Big{]},$ (7)
$\displaystyle\nabla_{\omega_{t+K-1}}\delta_{\omega_{t+K-1}}(t+l)=\gamma\phi_{\theta_{t+K-1}}(S_{t+l+1})-\phi_{\theta_{t+K-1}}(S_{t+l}).$
The intuition behind the critic parameter update is explained in Appendix A.7.
The actor-critic algorithm.
We combine parameter updates in Definitions 4.1 and 4.3 under the architecture illustrated in Figure 1. The full arrows depict the flow of state observations from the environment through the feature extractor and the actor’s network. The dashed arrows depict the flow of state observations from the environment through the feature extractor and the critic’s network.
$V_{\omega}(s)$$\pi_{\theta}(a|s)$$\mathcal{S}$$\pi_{\theta}$$\omega^{T}\phi_{\theta}$$\phi_{\theta}$
Figure 1: Actor-Critic Architecture: Actor and critic share features $\phi_{\theta}$ and use them separately to produce the current action and state-value, respectively. The gradient computation in 4.1 and 4.3 is executed independently with addition that the actor’s back-propagation is used to update the weights in the feature extractor as well.
5 Convergence analysis of the actor-critic algorithm
For a given matrix $M\in\mathbb{Z}^{n\times d}$ and a starting point $u\in\mathbb{Z}_{\geq 0}^{d}$, we will refer to the algorithm which computes a lattice basis of $M$ and applies A2C via architecture 1 as the “Actor-Critic Move Generator” (Algorithm 2). Its policy is used to explore the fiber through the MDP formulation in 3.2.
Theorem 5.1.
Assume that the parameter spaces $\Theta$ and $\Omega$ are closed and bounded. Given the finite fiber $\mathcal{F}_{M}$, bounded reward $R_{t}$, and parameter updates in Equations (4.1) and (4.3), the Actor-Critic Move Generator converges to the approximate optimal policy $\pi^{*}$ such that
$\displaystyle\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\pi}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}(S_{t},\pi(A_{t}|S_{t}),S_{t+1})\Big{]}.$
Proof intuition.
The proof of convergence of the Actor-Critic algorithm relies on the 2-timescale ordinary differential equation analysis pioneered by [10]. This method has been used to analyze various versions of the actor-critic method; see [9],[39], [11]. The tools that we use to show convergence rely on 6 technical conditions, all of which are stated and proved in Appendix A.8, pages A.8-A.10. Given two assumptions (4.1, and 4.2) on the compactness of the parameter space of the actor and critic, as well as a standard assumption on the ergodicity (A.1) of the Markov chains simulated in the Markov decision process from Definition 3.2, we show that our algorithm satisfies these 6 conditions and hence converges.
To address the computational complexity of determining a lattice basis for large $M$, note the following.
Corollary 5.1.
Subdividing the problem, applying the Actor-Critic Move Generator, and then embedding the resulting policy into the original problem, preserves optimality.
Exactly how problem subdivision is achieved is up to the user, and dependent on the particular problem and data structure. The next sections offers some examples.
6 Experiments
As discussed on page 3, data is a point $u\in\mathbb{Z}_{\geq 0}^{d}$, even if in statistical applications data arise in different formats. Here we consider data in the from of an adjacency matrix of an undirected graph on $N$ nodes, represented by a vertex-edge incidence vector in which $d$ is the number of node-pairs. We sample the fibers of the $\beta$-model for random graphs [12], with the design matrix $M\in\mathbb{Z}^{N\times d}$. Data generation and computing times are detailed in Appendix A.9.
Methodology. Since the Actor-Critic Move Generator is applied to subproblems, we can employ special algorithms to decompose graphs and then reconstruct them via Corollary 5.1. The only rule in the problem subdivision step is that the instances need to be small enough so that lattice bases of the null space can be computed, but also contain enough information to yield a conclusive answer to the proposed statistical test. For graph data, we use three different methods for initial data subdivision: $k$-core decomposition, bridge cuts, and induced subgraphs; this choice is governed by data sparsity. The pseudo code for the decomposition algorithm as well as Actor-Critic Move Generator, permutation and random sampling algorithms can be found in A.10.
Performance evaluation metrics. For the co-authorship dataset, we report the hypothesis test conclusion and compare it with known statistical results. For the randomly generated graphs, the Erdös-Rényi model is too simple to test, so instead of statistical conclusions, we report two other metrics: the number of discovered states in the fiber given different levels of sparsity, showing we can obtain an exchangeable fiber sample; and the number of discovered states in the fiber which the agent hasn’t learned on, looking to deduce how well the algorithm generalizes to unseen data.
Figure 2: (Left) A component of statistician coauthorship network with 2323 2323 nodes. (Right) Sampling distribution of the chi-square statistic on the fiber; sample size $50,000$. The observed value is in red.
Fiber Exploration
N $p$ $p=0.5$ $p=0.3$ $p=0.1$
1000 $6.9e{29}$ $1.8e{12}$ $3.6e{9}$
2000 $2.3e{29}$ $1.1e{11}$ $3.7e{7}$
3000 $1.9e{29}$ $3.4e{29}$ $16\displaystyle 4.8e$ 16
Table 1: Illustration of Actor-Critic Move Generator scaling, and the negligible effect of data sparsity on fiber exploration. Table entries are the number of discovered fiber points given a random initial vector drawn from the Erdös-Rényi random graph model $G(N,p)$.
Results. The three main takeaways from the simulation results, summarizing our evaluation of algorithm performance, scaling, and generalization, are as follows. (1) A model-based approach for analyzing a component of the coauthorship network, which we use as a benchmark, is offered in [29]. Our statistical conclusions are consistent with the known results: the $\beta$ model fits the data well, with the $p$-value 0.9957. The sampling distribution of the goodness-of-fit statistic in Figure 2 also resembles that of the benchmark, indicating a statistically good fiber sample. (2) For simulated data, we test the difficulty of sampling randomly generated graphs with different sparsity levels. Table 1, which reports the number of discovered fiber points using the Actor-Critic Move Generator for subproblems after applying Corollary 5.1, illustrates that the algorithm scales and applies to data with various levels of sparsity. We see that the algorithm is able to sample a very large number of fiber points, relative to the edge probability parameter $p$ and number of nodes $N$. (3) The final point is the test of how well the algorithm generalizes to a problem it has not seen before. Similar to [51], we train the agent on a sparse problem, and then use the generated moves on another–more dense–dataset and report the number of discovered states. We are interested in seeing how easy it is to apply this algorithm to another dataset without any learning involved. Figure 3 shows that the number of discovered states increases when we apply the learned moves to denser fibers.
Fibers of a 5050 5050-node graph $g\in\mathbb{Z}^{1225}$.
Fibers of a 10001000 10001000-node graph $g\in\mathbb{Z}^{499500}$.
Figure 3: We show the generalization properties of our move generation algorithm. Depicted is the plot of the $95\%$ confidence interval of discovered states with 300300 300300 generated moves applied to unseen fibers with increasing $L_{1}$ norm on the lower bound on the entry of the vector $b=Mu$.
7 Conclusion
Summary. We have demonstrated a novel approach for sampling lattice points in high-dimensional polytopes using reinforcement learning. Leveraging the power of parameterized policies and identifying substructures in the problem allows addresses both the dimensionality and sparsity bottlenecks that naturally arise in such problems. A goodness-of-fit test for sparse categorical data using an exchangeable sample constructed with our algorithm matches statistical conclusions in published work [29] for a popular network data set [27]. Robustness and generalization properties of our method are demonstrated on synthetic network data with varying sparsity levels. We have proved the asymptotic convergence of the algorithm.
Future work. We plan to expand the application of reinforcement learning to massive fiber sampling problems. Specifically, once the initial solution is split into sub-problems, a small part of the information is removed. To incorporate this information back into the sampler, one can recalibrate it by a double-sweep of the initial solution to produce overlapping subproblems. Initial tests indicate that such a strategy promises to ensure better fiber exploration through a sample with richer features.
8 Acknowledgements
This work is partially supported by the Simons Foundation’s Travel Support for Mathematicians Gift # 854770 and DOE/SC award number #1010629. Both authors are with Illinois Institute of Technology, where IG is a PhD student.
Appendix A Appendix
A.1 Detailed comparison to related work
[51] uses two key theoretical ingredients for their game. First, using a powerful polytope universality result, they show that every integer feasibility problem is linearly equivalent to a 33 33-dimensional problem where certain entries of the solution vector must be zero. This low-dimensional problem’s feasibility and an initial solution are easily determined. Then, they propose to use RL to traverse the space of all of its solutions and reward the agent for discovering the states that have zeros in the entries prescribed by the linear embedding. Thus, the winning state will be a feasible solution to the original problem. The second ingredient are the actions used to traverse the solution space of the 33 33-dimensional problem, known in computational algebra as a Gröbner basis [42]. One of the main computational challenges in this setting is the fact that the low-dimensional embedding is too large to explore. Hence the application of RL is a logical step forward due to its ability to scale with the problem size, as well as produce an approximate optimal answer to the problem, without searching through the entire solution space. In terms of RL, they use a twin delayed deep deterministic policy gradient algorithm (DDPG): an off-policy RL algorithm that uses two action-value networks $Q(\cdot,\cdot\ ;\theta)$ and $Q(\cdot,\cdot\ ;\theta^{\prime})$ to encode the latest optimal policy of the agent. The two $Q$ networks help with prevention of overestimation of the Q-function during training. Furthermore, because the targets in the learning regime are Bellman updates given by the sum of the current reward and the discounted future reward, under the constantly updated policy, they are in a sense non-stationary. Hence, occasionally freezing the weights of $Q(\cdot,\cdot\ ;\theta^{\prime})$ and using it for the Bellman target computation and using $Q(\cdot,\cdot\ ;\theta)$ for storing the optimal policy helps with minimization of the loss function during training. However, such an approach requires one to keep a large enough memory buffer, as well as a burn-in sample, in order to train both networks. During training, the algorithm uses samples from both the memory buffer and the burn-in sample to de-correlate the underlying dependency of the sample with respect to the current policy $\pi$. The reason is that one desires an i.i.d. sample during the network training stage to obtain an unbiased estimator of the action-value function. They demonstrate the method on the equivalent two-dimensional problem.
In our algorithm, we do not use a Gröbner basis, but rather a lattice basis, a subset order of magnitude smaller (a smallest possible Markov sub-basis [13, Lemma 4.2]). We consider the general fiber sampling problem, rather than the integer feasibility problem, and note that we can apply our algorithm to the problem they are solving, by using modifying our reward function $R_{t}$ to include a cost minimization aspect to it. Another key difference between our work and [51] is the RL algorithm that we implement. The off-policy approach will likely fail in the case when one is working with sparse fibers, where obtaining a burn-in sample using some greedy algorithm has proved a nontrivial task. For example, if we naively apply a random walk algorithm to a sparse initial solution, where each move is sampled from a multinomial distribution, more often than not the random walk algorithm cannot produce a single feasible point besides the initial one (for a tiny but sparse example with low move acceptance probabilities, see [2, Figure 1]). Unlike the twin delayed DDPG, we do not require a burn-in sample of observations $(s_{t},a_{t},r_{t},s_{t+1})_{i},\ i\in[N]$, as we are using an on-line training method. This approach relies on training the policy and state-value neural networks on the sample collected along the current trajectory $(s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},\dots)$ consisting of states, actions and corresponding rewards. Furthermore, because we are deconstructing the initial solution to a collection of sub-problems before training, and using only a subset of lattice basis vectors per sub-problem, we are able to sample sparse fibers relatively effectively, restricting the magnitude of each move. Finally, we note that as in [51, §6.5], our algorithm generalizes well when perturbing the fiber by increasing the marginal vector $b$ (see Figure 3). Such a generalization is extremely helpful when working with limited computation resources, and it minimizes training times of the neural networks.
A.2 Ergodicity of the fiber sampling MDP
The next assumption ensures that the Markov chains obtained by following any policy $\pi$ have a stationary distributions, through which we can define loss functions in the expectation.
Assumption A.1.
Markov chains $\{S_{t}\}$ and $\{S_{t},A_{t}\}$ are irreducible and aperiodic with a unique stationary probabilities $\mu(s)$ and $\eta(s,a)=\mu(s)\pi(a|s)$.
In our case the above assumption is satisfied due to the finite state space $\mathcal{S}$: in the problems with which we work, $P_{A}(b)$ is a polytope, which means it is bounded, and therefore the fiber is finite. Any state can be reached from any other state, in finite time. And moreover, aperiodic property is satisfied because we are able to produce a zero action $A_{t}=\textbf{0}$. [24] assume uniqueness of measure, and we follow the same setup.
In general, the linear system $Mx=b$ may define a polyhedral system which is unbounded, in which case the state space would be infinite. Our algorithm assumes a polytope.
A.3 Cumulative discounted reward
Let $G_{t}$ denote the cumulative discounted reward starting at time $t$ and write
$\displaystyle G_{t}:=\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}.$
Because we are sampling a finite fiber $\mathcal{F}_{A}$, the state space is bounded, i.e. there exists $B_{\mathcal{S}}$ such that for all $s\in\mathcal{S}$, it follows that $|s|\leq B_{\mathcal{S}}$. Furthermore, the action space $\mathcal{A}$ is also bounded by design, mentioned in Remark 3.1. Therefore, we have that for all $t$, reward $R_{t+1}$ is also bounded, i.e. for all $(S_{t},A_{t},S_{t+1})$ triple, there exists a constant $B_{R}$ such that
$\displaystyle|R_{t+1}|\leq B_{R}.$
Then, it immediately follows that
$\displaystyle G_{t}\leq\frac{1}{1-\gamma}B_{R},\ \gamma\in(0,1)$
by the convergence of a geometric series. Let $V_{\pi}$ denote the state-value function under policy $\pi$, starting from state $S_{t}=s$ at time $t$, given by
$\displaystyle V_{\pi}(s):=\mathbb{E}_{\tau\sim\pi}\Big{[}G_{t}\Big{|}\ S_{t}=s\Big{]}$
where $\tau$ denotes the trajectory sampled using policy $\pi$. Similarly as above, it follows that
$\displaystyle V_{\pi}(s)\leq\frac{1}{1-\gamma}B_{R},\ \gamma\in(0,1).$
Furthermore, denote the action-value function by $Q_{\pi}(s,a)$ under policy $\pi$ and write
$\displaystyle Q_{\pi}(s,a):=\mathbb{E}_{\tau\sim\pi}\Big{[}G_{t}\Big{|}\ S_{t}=s,A_{t}=a\Big{]}.$
Again it follows that
$\displaystyle Q_{\pi}(s,a)\leq\frac{1}{1-\gamma}B_{R},\ \gamma\in(0,1).$
A.4 Bounded policy gradient
Proof of Claim 4.1.
Part (a) of the claim follows immediately from ergodicity, stated in Assumption A.1. Let us prove part (b). First note that for any parametarization $\theta$, the policy $\pi(a|s;\theta)$ is bounded for all $s,a$ pairs:
$\displaystyle 0<\pi(a|s;\theta)\leq 1,\ \leavevmode\nobreak\ \text{for all }\theta,s,a.$
Furthermore, the map $\theta\mapsto\pi_{\theta}(a|s)$ is twice continuously differentiable if we choose a smooth enough activation functions of the network $h$; for example, one can pick sigmoid, tanh, ELU, etc.). Next, fix $s$ and $a$. Taking the natural log of the policy gives:
$\displaystyle\ln\pi(a|s;\theta)=h(a,s;\ \theta)-\ln\sum_{s^{\prime}}e^{h(a,s^{\prime};\ \theta)}$
and taking the gradient, we obtain:
$\displaystyle\nabla_{\theta}\ln\pi(a|s;\theta)$ $\displaystyle=\nabla_{\theta}h(a,s;\ \theta)-\frac{\sum_{s^{\prime}}e^{h_{\theta}}\nabla_{\theta}h_{\theta}}{\sum_{s^{\prime}}e^{h_{\theta}}}$
$\displaystyle=\nabla_{\theta}h(a,s;\ \theta)-\sum_{s^{\prime}}\pi(a|s^{\prime};\theta)\nabla_{\theta}h_{\theta}$
$\displaystyle=\nabla_{\theta}h(a,s;\ \theta)-\mathbb{E}_{\pi_{\theta}}\Big{[}\nabla_{\theta}h_{\theta}\Big{]}.$
Let $f(\theta):=\nabla_{\theta}\ln\pi(a|s;\theta)$. Due to Assumption 4.1, it follows that $f(\theta)$ is bounded. Taking a derivative of $f$ with respect to $\theta$, it follows that
$\displaystyle\frac{d}{d\theta}f(\theta)=\nabla_{\theta}^{2}\ln\pi(a|s;\theta)=\nabla_{\theta}^{2}h(a,s;\ \theta)-\sum_{s^{\prime}}\nabla_{\theta}\pi(a|s^{\prime};\theta)\nabla_{\theta}h_{\theta}+\pi(a|s^{\prime};\theta)\nabla_{\theta}^{2}h_{\theta}.$
Again, due to Assumption 4.1, the derivative $\frac{d}{d\theta}f(\theta)$ is bounded as well. ∎
A.5 Derivation of the actor parameter update
The objective function $L_{a}(\theta)$ of the actor, under policy $\pi_{\theta}$ parametrized by $\theta$,
$L_{a}(\theta):=V_{\pi_{\theta}}(s_{0}).$
This expression is exactly the state-value function starting from the initial state $S_{0}=s_{0}$. Next, consider the classical theorem which computes the gradient of the state-value function.
Theorem A.1 (Policy Gradient Theorem [44]).
Given the state-value function $V_{\pi_{\theta}}$ under parametrized policy $\pi_{\theta}$, define the gradient objective $J(\theta)$ as
$\displaystyle L_{a}(\theta)=V_{\pi_{\theta}}(s_{0}),$
where $\mu_{\pi_{\theta}}$ is the stationary distribution of states $s$ under the Markov chain induced by policy $\pi_{\theta}$. Then, the gradient of the objective $L_{a}(\theta)$ with respect to to parameter $\theta$ is given by
$\displaystyle\nabla_{\theta}L_{a}(\theta)=\nabla_{\theta}V_{\pi_{\theta}}(s_{0})=\mathbb{E}_{\pi_{\theta}}\Big{[}Q_{\pi_{\theta}}(S_{t},A_{t})\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}.$
Proof.
Note that the following relationship holds, $Q(s,a)=R_{t+1}(s,a)+\gamma V(S_{t+1})$, because $R_{t+1}$ does not depend on $S_{t+1}$ hence it can be moved outside of the expectation. Then, given an arbitrary observed state $S_{t}=s$ at time $t$, we can write:
$\displaystyle\nabla_{\theta}V_{\pi_{\theta}}(s)$ $\displaystyle=\nabla_{\theta}\mathbb{E}_{\pi_{\theta}}\Big{[}G_{t}\Big{|}\ S_{t}=s,\pi_{\theta}\Big{]}=\nabla_{\theta}\mathbb{E}_{\pi_{\theta}}\Big{[}R_{t+1}(s,A_{t})+\gamma G_{t+1}\Big{|}\ S_{t}=s,\pi_{\theta}\Big{]}$
$\displaystyle=\nabla_{\theta}\mathbb{E}_{\pi_{\theta}}\Big{[}R_{t+1}(s,A_{t})+\gamma V_{\pi_{\theta}}(S_{t+1})\Big{|}\ S_{t}=s,\pi_{\theta}\Big{]}=\nabla_{\theta}\sum_{a}\pi_{\theta}(a|s)\Big{[}R_{t+1}(s,a)+\gamma V_{\pi_{\theta}}(S_{t+1})\Big{]}$
$\displaystyle=\nabla_{\theta}\sum_{a}\pi_{\theta}(a|s)Q_{\pi_{\theta}}(a,s)=\sum_{a}\Big{[}\nabla_{\theta}\pi_{\theta}(a|s)Q_{\pi_{\theta}}(a,s)+\pi_{\theta}(a|s)\nabla_{\theta}Q_{\pi_{\theta}}(a,s)\Big{]}$
$\displaystyle=\sum_{a}\Big{[}\nabla_{\theta}\pi_{\theta}(a|s)Q_{\pi_{\theta}}(a,s)+\pi_{\theta}(a|s)\nabla_{\theta}\sum_{s^{\prime},r}p(s^{\prime},r|s,a)(R_{t+2}+\gamma V_{\pi_{\theta}}(s^{\prime}))\Big{]}$
$\displaystyle=\sum_{a}\Big{[}\nabla_{\theta}\pi_{\theta}(a|s)Q_{\pi_{\theta}}(a,s)+\pi_{\theta}(a|s)\sum_{s^{\prime}}p(s^{\prime}|s,a)\gamma\nabla_{\theta}V_{\pi_{\theta}}(s^{\prime})\Big{]}$
$\displaystyle=\sum_{a}\bigg{[}\nabla_{\theta}\pi_{\theta}(a|s)Q_{\pi_{\theta}}(a,s)+\pi_{\theta}(a|s)\sum_{s^{\prime}}p(s^{\prime}|s,a)\cdot$
$\displaystyle\gamma\sum_{a^{\prime}}\Big{[}\nabla_{\theta}\pi_{\theta}(a^{\prime}|s^{\prime})Q_{\pi_{\theta}}(a^{\prime},s^{\prime})+\pi_{\theta}(a^{\prime}|s^{\prime})\sum_{s^{\prime\prime}}p(s^{\prime\prime}|s^{\prime},a^{\prime})\gamma\nabla_{\theta}V_{\pi_{\theta}}(s^{\prime\prime})\Big{]}\bigg{]}$
$\displaystyle=\sum_{x\in\mathcal{S}}\sum_{k=0}^{\infty}\gamma^{k}Pr(s\rightarrow x,k,\pi_{\theta})\sum_{a}\nabla_{\theta}\pi_{\theta}(a|x)Q_{\pi_{\theta}}(a,x).$
Here $Pr(s\rightarrow x,k,\pi_{\theta})$ is the probability of being in state $x$ after $k$ steps under policy $\pi_{\theta}$ and starting from state $s$. Note that the discounting factors $\gamma,\gamma^{2},...$ have been absorbed into the probability. It follows then that
$\displaystyle\nabla_{\theta}L_{a}(\theta)$ $\displaystyle=\nabla_{\theta}V_{\pi_{\theta}}(s_{0})=\sum_{s}\sum_{k=0}^{\infty}Pr(s_{0}\rightarrow s,k,\pi_{\theta})\sum_{a}\nabla_{\theta}\pi_{\theta}(a|s)Q_{\pi_{\theta}}(a,s)$
$\displaystyle=\sum_{s}\eta(s)\sum_{a}\nabla_{\theta}\pi_{\theta}(a|s)Q_{\pi_{\theta}}(a,s)=\sum_{s^{\prime}}\eta(s^{\prime})\sum_{s}\frac{\eta(s)}{\sum_{s^{\prime}}\eta(s^{\prime})}\sum_{a}\nabla_{\theta}\pi_{\theta}(a|s)Q_{\pi_{\theta}}(a,s)$
$\displaystyle=\sum_{s^{\prime}}\eta(s^{\prime})\sum_{s}\mu_{\pi_{\theta}}(s)\sum_{a}\nabla_{\theta}\pi_{\theta}(a|s)Q_{\pi_{\theta}}(a,s)$
$\displaystyle\propto\sum_{s}\mu_{\pi_{\theta}}(s)\sum_{a}\nabla_{\theta}\pi_{\theta}(a|s)Q_{\pi_{\theta}}(a,s),$
where $\mu_{\pi_{\theta}}(s)$ is the on-policy, stationary state distribution under policy $\pi_{\theta}$. In the case of a continuing task, the proportionality constant is 11 11, therefore
$\displaystyle\nabla_{\theta}L_{a}(\theta)$ $\displaystyle=\sum_{s}\mu_{\pi_{\theta}}(s)\sum_{a}\nabla_{\theta}\pi_{\theta}(a|s)Q_{\pi_{\theta}}(s,a).$
We can further manipulate the expression above to get a gradient computation found in the REINFORCE algorithm [50]:
$\displaystyle\nabla_{\theta}L_{a}(\theta)$ $\displaystyle=\sum_{s}\mu_{\pi_{\theta}}(s)\sum_{a}\nabla_{\theta}\pi_{\theta}(a|s)Q_{\pi_{\theta}}(s,a)$
$\displaystyle=\mathbb{E}_{\pi_{\theta}}\Big{[}\sum_{a}\nabla_{\theta}\pi_{\theta}(a|S_{t})Q_{\pi_{\theta}}(S_{t},a)\Big{]}$
$\displaystyle=\mathbb{E}_{\pi_{\theta}}\Big{[}\sum_{a}\pi_{\theta}(a|S_{t})Q_{\pi_{\theta}}(S_{t},a)\frac{\nabla_{\theta}\pi_{\theta}(a|S_{t})}{\pi_{\theta}(a|S_{t})}\Big{]}$
$\displaystyle=\mathbb{E}_{\pi_{\theta}}\Big{[}Q_{\pi_{\theta}}(S_{t},A_{t})\frac{\nabla_{\theta}\pi_{\theta}(A_{t}|S_{t})}{\pi_{\theta}(A_{t}|S_{t})}\Big{]}$
$\displaystyle=\mathbb{E}_{\pi_{\theta}}\Big{[}Q_{\pi_{\theta}}(S_{t},A_{t})\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}.$
We can modify the gradient computation above in a way that the value of the gradient remains the same, while the variance of an estimator of this gradient is reduced. We accomplish this by introducing a baseline function, denoted $b(\cdot)$. The following Lemma A.1 shows that this modification indeed does not change the policy gradient.
Lemma A.1.
Given the baseline function $b(S_{t})$, we can write the gradient of the state-value function in the form of
$\displaystyle\nabla_{\theta}L_{a}(\theta)$ $\displaystyle=\mathbb{E}_{\pi_{\theta}}\Big{[}\big{(}Q_{\pi_{\theta}}(S_{t},A_{t})-b(S_{t})\big{)}\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}.$
Proof.
We want to show that
$\displaystyle\mathbb{E}_{\pi_{\theta}}\Big{[}b(S_{t})\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}=0,$
from which the result will follow. We note that one can rewrite the expectation as:
$\displaystyle\mathbb{E}_{\pi_{\theta}}\Big{[}b(S_{t})\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}$ $\displaystyle=\sum_{S_{t}=s}\mu_{\pi_{\theta}}(s)\sum_{A_{t}=a}\nabla_{\theta}\pi_{\theta}(a|s)b(s)$
$\displaystyle=\sum_{S_{t}=s}\mu_{\pi_{\theta}}(s)b(s)\nabla_{\theta}\sum_{A_{t}=a}\pi_{\theta}(a|s)$
$\displaystyle=\sum_{S_{t}=s}\mu_{\pi_{\theta}}(s)b(s)\nabla_{\theta}1=0.$
Thus, it follows that
$\displaystyle\nabla_{\theta}L_{a}(\theta)$ $\displaystyle=\mathbb{E}_{\pi_{\theta}}\Big{[}\big{(}Q_{\pi_{\theta}}(S_{t},A_{t})-b(S_{t})\big{)}\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}=\nabla_{\theta}L_{a}(\theta)-0=\nabla_{\theta}L_{a}(\theta).$
Therefore, as long as the baseline function does not depend on the action $A_{t}=a$, we can absorb it inside of the gradient computation. For example, a popular choice is $b(S_{t})=V_{\pi_{\theta}}(S_{t})$. Then, we get
$\displaystyle A_{\theta}(S_{t},A_{t})=Q_{\pi_{\theta}}(S_{t},A_{t})-V_{\pi_{\theta}}(S_{t}),$
which is called the advantage function [3]. Then the gradient is written as
$\displaystyle\nabla_{\theta}L_{a}(\theta)=\mathbb{E}_{\pi_{\theta}}\Big{[}A_{\theta}(S_{t},A_{t})\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}.$ (8)
Notice that we can approximate the gradient in (8) by defining an estimator
$\displaystyle\widehat{\nabla_{\theta}L_{a}(\theta)}:=\frac{1}{K}\sum_{l=0}^{K-1}\delta_{\theta}(l)\nabla_{\theta}\ln\pi_{\theta}(A_{l}|S_{l}),$
where
$\displaystyle\delta_{\theta}(l):=R_{l+1}(S_{l},A_{l})+\gamma V_{\pi_{\theta}}(S_{l}+A_{l})-V_{\pi_{\theta}}(S_{l})$
is called a temporal difference (TD) [43]. Note that $S_{l+1}=S_{l}+A_{l}$ due to the environment dynamics defined in 3.2. We can go a step further and generalize the estimate above.
Definition A.1 (Generalized Advantage Estimate (GAE) [40]).
Define the advantage estimate in the following way
$\displaystyle\hat{A}^{(1)}(t)$ $\displaystyle=R_{t+1}+\gamma V(S_{t+1})-V(S_{t})=\delta(t)$
$\displaystyle\hat{A}^{(2)}(t)$ $\displaystyle=R_{t+1}+\gamma R_{t+2}+\gamma^{2}V(S_{t+2})-V(S_{t})=\delta(t)+\gamma(\delta(t+1))$
[…]
$\displaystyle\hat{A}^{(k)}(t)$ $\displaystyle=\sum_{l=0}^{k}\gamma^{l}\delta(t+l)$
[…]
$\displaystyle\hat{A}^{(\infty)}(t)$ $\displaystyle=\sum_{l=0}^{\infty}\gamma^{l}\delta(t+l)=-V(S_{t})+\sum_{l=0}^{\infty}\gamma^{l}R_{t+1+l}.$
Then, for $\lambda\in[0,1]$ and policy $\pi_{\theta}$, define the Generalized Advantage Estimate as
$\displaystyle\widehat{A_{GAE}}(t)=(1-\lambda)\sum_{k=0}^{\infty}\lambda^{k}\hat{A}^{(k)}(t)=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta(t+l).$
Note that the GAE is a biased estimator of the advantage function $A_{\theta}(S_{t},A_{t})=Q_{\pi_{\theta}}(S_{t},A_{t})-V_{\pi_{\theta}}(S_{t})$, in which the bias comes from the discount factor $\lambda$. It decomposes into an estimator of $Q_{\pi_{\theta}}(S_{t},A_{t})$ and the baseline function $b(S_{t})$. We thus define a generalized advantage estimator policy gradient as:
$\displaystyle\nabla_{\theta}L_{GAE}(\theta):=\mathbb{E}_{\pi_{\theta}}\Big{[}\widehat{A_{GAE}}(t)\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}.$
Furthermore, note that because the TDs are bounded, due to the state-value function and instantaneous rewards being bounded, the $\widehat{A_{GAE}}$ is bounded as well for all $t$. From the previous subsection we know that $\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})$ is bounded as well and thus, by the Fubini-Tonelli theorem, it follows that
$\displaystyle\nabla_{\theta}L_{GAE}(\theta)=\mathbb{E}_{\pi_{\theta}}\Big{[}\widehat{A_{GAE}}(t)\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\mathbb{E}_{\pi_{\theta}}\Big{[}\delta(t+l)\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}.$
Then we have that
$\displaystyle\nabla_{\theta}L_{a}(\theta)=\mathbb{E}_{\pi_{\theta}}\Big{[}A_{\theta}(S_{t},A_{t})\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}\approx\nabla_{\theta}L_{GAE}(\theta)$
where equality holds when $\lambda=1$. Note that the TD terms $\delta(k+l)$ are exponentially discounted and thus terms further in the future contribute significantly less to the value of the infinite sum. Hence, we can truncate the sum up to some step $K-1$ such that
$\displaystyle g_{a}(\theta)$ $\displaystyle:=\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\mathbb{E}_{\pi_{\theta}}\Big{[}\delta(t+l)\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}$ (9)
$\displaystyle=\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\Big{(}\mathbb{E}_{\pi_{\theta}}\Big{[}R_{t+1}(S_{t},A_{t})\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}$
$\displaystyle+\mathbb{E}_{\pi_{\theta}}\Big{[}\gamma V_{\pi_{\theta}}(S_{t+1})\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}$
$\displaystyle-\mathbb{E}_{\pi_{\theta}}\Big{[}V_{\pi_{\theta}}(S_{t})\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}\Big{)}.$
All the above computations have been derived under the assumption that we have an exact expression of the state-value function $V_{\pi_{\theta}}$. However, in practice, this is usually not the case and we can only work with an approximation of such function under some parametarization $\omega$. In what follows, we show that such parametarization does not pose any issue in the gradient estimate. We follow the work of [31] and [47] and define, for any $\theta$, an inner product $\langle\cdot,\cdot\rangle_{\theta}$ of two real-valued functions $F_{1},F_{2}$, on $\mathcal{S}$ and $\mathcal{S}\times\mathcal{A}$, viewed as vectors in $\mathbb{Z}^{|\mathcal{S}|}$ and $\mathbb{Z}^{|\mathcal{S}||\mathcal{A}|}$, by
$\displaystyle\langle F_{1},F_{2}\rangle_{\theta}:=\sum_{s}\mu_{\theta}(s)\sum_{a}\pi_{\theta}(a|s)F_{1}(s)F_{2}(s,a)=\sum_{s}\mu_{\theta}(s)F_{1}(s)\sum_{a}\pi_{\theta}(a|s)F_{2}(s,a).$
The functions $F_{1}$ and $F_{2}$ are viewed as vectors in a way that each state $s$ represents a single coordinate. If we only consider partial derivative w.r.t parameter components $\theta^{i}$ we can write equation (9) as
$\displaystyle\frac{\partial}{\partial\theta^{i}}g_{a}(\theta)=\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\langle\delta(t+l),\psi_{\theta}^{i}\rangle_{\theta},\ \leavevmode\nobreak\ i=1,\dots,n,$
where $\psi_{\theta}^{i}$ denotes the $i-th$ vector component of
$\displaystyle\psi_{\theta}(S_{t}):=\sum_{a}\pi_{\theta}(a|S_{t})\nabla_{\theta}\ln\pi_{\theta}(a|S_{t}).$
To clarify the last sentence, consider the object $\psi_{\theta}(S_{t})$, which is already a vector of a fixed state $S_{t}$, due to the gradient term inside it. However, in the view of the inner product $\langle\cdot,\cdot\rangle_{\theta}$ each component is a different state $s$ and not the partial derivative of the gradient itself. Thus, we are facing a vector with components being vectors as well. The outer vector is comprised of state components $S_{t}=s$ and the inner vectors are comprised of the gradient with respect to $\theta$. Hence, if we only consider the component of the gradient indexed by the partial derivative $\frac{\partial}{\partial\theta^{i}}$, we get vectors $\psi_{\theta}^{i}$ in $\mathbb{Z}^{|\mathcal{S}|}$. Furthermore, define a projection $\Pi_{\theta}:\mathbb{Z}^{|\mathcal{S}|}\mapsto\Psi_{\theta}$ as
$\displaystyle\Pi_{\theta}F:=\arg\min_{\hat{F}\in\Psi_{\theta}}||F-\hat{F}||_{\theta}$
[…]
$\displaystyle\langle\Pi_{\theta}V_{\pi_{\theta}}-V_{\pi_{\theta}},\psi_{\theta}^{i}\rangle_{\theta}=0$
$\displaystyle\langle\Pi_{\theta}V_{\pi_{\theta}},\psi_{\theta}^{i}\rangle_{\theta}=\langle V_{\pi_{\theta}},\psi_{\theta}^{i}\rangle_{\theta}$
which shows that it suffices to learn the projection of $V_{\pi_{\theta}}$ onto $\Psi_{\theta}$ in order to compute the approximation of the gradient $g_{a}(\theta)$. So, let $V_{\omega}$ be a projection of the state-value function onto the subspace spanned by $\Psi_{\theta}$, and let
$\displaystyle\delta_{\omega}(t):=R_{t+1}+\gamma V_{\omega}(t+1)-V_{\omega}(t).$
Then we write the gradient approximation of the exact gradient as
$\displaystyle g_{a}(\theta,\omega):=\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\mathbb{E}_{\pi_{\theta}}\Big{[}\delta_{\omega}(t+l)\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})\Big{]}.$ (10)
Furthermore, because we are collecting data in a on-line fashion, we are estimating the gradient along a truncated trajectory $\tau_{t:t+K}$. Hence, the gradient estimator can be computed as
$\displaystyle\hat{g_{a}}(\theta,\omega):=\frac{1}{K}\sum_{k=0}^{K-1}\nabla_{\theta}\ln\pi_{\theta}(A_{k}|S_{k})\sum_{l=k}^{K-1}(\gamma\lambda)^{l-k}\delta_{\omega}(t+l).$ (11)
These detailed considerations allow us to define the actor parameter update as in Definition 4.1.
A.6 Feature vector bounds
In Section 4.1, we noted that, not having access to the analytical form of the state-value function $V$, we parametrize it:
$V_{\omega}(S_{t}):=\phi_{\theta}(S_{t})^{T}\omega\approx V_{\theta}.$
This parametrization can be written in matrix form as follows:
$\displaystyle V_{\omega}(S_{t})=\begin{bmatrix}&\phi_{\theta}(S_{1})\\ &\vdots\\ &\phi_{\theta}(S_{n})\end{bmatrix}\omega=\begin{bmatrix}&\phi_{1,\theta}(S_{1})&\dots&\phi_{n,\theta}(S_{1})\\ &\vdots\\ &\phi_{1,\theta}(S_{n})&\dots&\phi_{n,\theta}(S_{n})\end{bmatrix}\omega=\Phi\omega.$
Claim A.1.
Adopting the notation above, the following conditions are satisfied:
(a) For every $s\in\mathcal{S}$, the mapping $\theta\mapsto\phi_{\theta}$ is continuous, bounded, and differentiable.
(b) The span of vectors $\phi_{\theta}^{j}$, $j=1,\dots,m$ in $\mathbb{Z}^{|\mathcal{S}|}$, denoted by $\Phi_{\theta}$ contains $\Psi_{\theta}$.
Proof of Claim A.1.
Part $(a)$ of the claim follows by assumption 4.1 and Claim 4.1.
Part $(b)$ is a technicality needed to ensure that the neural network is well posed. Essentially, it says the following. Given the feature vector space $\Phi_{\theta}$, the policy network maps the vector space to another vector space defined implicitly by $\pi_{\theta}(a|\phi_{\theta}(S_{t}))$. Taking the logarithm and then back-propagating with a gradient $\nabla_{\theta}$, we again end up in the feature space $\Phi_{\theta}$. Thus the assumption ensures that taking the gradient of the network, we will end back in the feature space. Such an assumption is naturally satisfied with a proper choice of activation functions. It also ensures that the projector $\Pi_{\theta}$ does not become ill-conditioned. ∎
A.7 Critic minimization objective and parameter update
As mentioned on page 4.1, one of the most basic but well established minimization objectives minimizes the difference between collected rewards along a truncated trajectory $\tau_{t:t+K}$ and the parametrized state-value function $V_{\omega}$. Focusing on the temporal difference in this objective, we replace it by a truncated version of the generalized advantage estimators.
In mathematical detail, consider the difference between collected rewards and state-value function:
$\displaystyle\min_{\omega}\Big{|}\Big{|}\Big{(}\sum_{k=0}^{K-1}\gamma^{k}R_{t+1+k}-V_{\omega}\Big{)}^{2}\Big{|}\Big{|}_{\theta}^{2},$ (12)
We employ a semi-gradient iterative scheme (cf. Section 9.3 in [44]) which uses bootstrapping for minimization, bootstrapping being necessary because it uses the current parameter $\omega$. The usual target in the minimization objective is the temporal difference $\delta_{\omega}$, as follows:
$\displaystyle\min_{\omega}\Big{|}\Big{|}\Big{(}R_{t+1}+\gamma V_{\omega}-V_{\omega}\Big{)}^{2}\Big{|}\Big{|}_{\theta}^{2}=\min_{\omega}\Big{|}\Big{|}\delta_{\omega}^{2}\Big{|}\Big{|}_{\theta}^{2}.$ (13)
However, we go a step further and use a truncated and parameterized version of $\widehat{A_{GAE}}(t)$ , denoted by $\widehat{A_{GAE,\omega}}(t)$ (cf. Definition A.1):
$\displaystyle\widehat{A_{GAE,\omega}}(t):=\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega}(t+l).$
Now, consider the definition of critic parameter update in 4.3. Assume that the system is in time $t+K-1$ and take the gradient of minimization objective (6) with respect to the parameter $\omega_{t+K-1}$:
$\displaystyle g_{c}(\theta_{t+K-1},\omega_{t+K-1})$ $\displaystyle=\nabla_{\omega_{t+K-1}}\Big{|}\Big{|}\frac{1}{2}\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t+K-1}}(t+l)^{2}\Big{|}\Big{|}_{\theta_{t+K-1}}^{2}$
$\displaystyle=\Big{|}\Big{|}\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t+K-1}}(t+l)\nabla_{\omega_{t+K-1}}\delta_{\omega_{t+K-1}}(t+l)\Big{|}\Big{|}_{\theta_{t+K-1}}^{2}.$ (14)
Approximating the norm with an observed sample of size $K$, we define the following gradient estimator at time $t+K-1$:
$\displaystyle\hat{g_{c}}(\theta_{t+K-1},\omega_{t+K-1}):=\frac{1}{K}\sum_{k=0}^{K-1}\Big{[}\sum_{l=k}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t+K-1}}(t+l)\nabla_{\omega_{t+K-1}}\delta_{\omega_{t+K-1}}(t+l)\Big{]}.$ (15)
A.8 Convergence: proof of Theorem 5.1
We aim to setup our algorithm in the framework presented by [10] and apply their convergence theorem. In the most general case, we are presented with the following stochastic iterative scheme
$\displaystyle x_{n+1}$ $\displaystyle=x_{n}+a(n)[h(x_{n},y_{n})+M_{n+1}^{(1)}]$ (16)
$\displaystyle y_{n+1}$ $\displaystyle=y_{n}+b(n)[g(x_{n},y_{n})+M_{n+1}^{(2)}]$ (17)
where $a(n)$ and $b(n)$ are step-sizes for the iterates, $h$ and $g$ are Lipschitz continues functions of the parameters at hand, and $M_{n+1}^{(i)}$ is a martingale difference with bounded second moments for $i=1,2$. The difficulty of showing convergence of such a scheme is that both equations above are coupled with respect to to the parameters they are updating. One potential approach would be to fix one parameter and converge the second, and then converge the first parameter as well. Although somewhat viable, the approximate convergence times for such approach would be much larger than what the o.d.e. method proposes. Namely, in this method the step-sizes $a(n)$ and $b(n)$ are chosen in such a way that one parameter is updated with a much slower rate than the other. In that case, the slower iterate (parameter) becomes quasi-stationary from the perspective of the fast iterate. By picking such viable step-sizes, one can show that the stochastic iterative scheme given by (16) and (17) converges to a local optimum $(x^{*},y^{*})$. The idea behind introducing ordinary differential equations in the analysis of the above scheme is that one can show that both iterates in (16) and (17) asymptotically follow the trajectories of certain o.d.e.’s. If these o.d.e.’s have an attractor (an equilibrium point), then the scheme converges to the same point.
The proof heavily leverages the setup presented above as well as the derivations in [31]. Before presenting the proof, we need to re-cast our algorithm in the form of Equtions (16) and (17) and check whether it satisfies the conditions presented above.
To this end, let $\theta_{t}$ and $\omega_{t}$ denote the actor and critic parameters when the system is in time $t$ respectfully. Our goal is to show that the parameters $(\theta_{t},\omega_{t})\longrightarrow(\theta^{*},\omega^{*})$ converge to a global optimum $(\theta^{*},\omega^{*})$. We do this by showing that iterates in Equations (5) and (7) satisfy conditions A.1, A.2, A.3, A.4, A.5, A.6, and then we apply Lemma A.2 and Theorem A.2.
We first put Equations (5) and (7) in the correct form. Let
$\displaystyle M_{t+1}^{(c)}=\hat{g_{c}}(\theta_{t},\omega_{t})-g_{c}(\theta_{t},\omega_{t}),$ (18)
$\displaystyle M_{t+1}^{(a)}=\hat{g_{a}}(\theta_{t},\omega_{t})-g_{a}(\theta_{t},\omega_{t}).$ (19)
Then the iterates above assume the following form
$\displaystyle\omega_{t+1}$ $\displaystyle=\omega_{t}+\beta(t)[g_{c}(\theta_{t},\omega_{t})+M_{t+1}^{(c)}]$ (20)
$\displaystyle\theta_{t+1}$ $\displaystyle=\theta_{t}+\alpha(t)[g_{a}(\theta_{t},\omega_{t})+M_{t+1}^{(a)}]$ (21)
where equations (20),(21) are now in the form of (16) and (17), respectively. In order for the Theorem 1 in [10] (also found in [11]) to be satisfied, we need to show that our algorithm satisfies the following conditions.
Condition A.1.
Functions $h:\mathbb{R}^{d+k}\longrightarrow\mathbb{R}^{d}$ and $g:\mathbb{R}^{d+k}\longrightarrow\mathbb{R}^{k}$ are continues and Lipschitz with respect to both input parameters.
Condition A.2.
Objects $M_{n+1}^{(i)}$, $i=1,2$ are martingale differences with bounded second moments, i.e.
$\displaystyle\mathbb{E}[M_{n+1}^{(i)}|\mathcal{I}_{n}]=0,\ \leavevmode\nobreak\ i=1,2$
where $\mathcal{I}_{n}=\sigma(x_{m},y_{m},M_{m}^{(1)},M_{m}^{(2)},m\leq n)$, $\mathcal{I}_{0}\subset\dots\subset\mathcal{I}_{n}\subset\mathcal{I}_{n+1}\subset\dots\subset\mathcal{I}$ is the filtration generated by information at subsequent times, consisted of prameters $x_{n},y_{n}$ and martingale differences $M_{m}^{(i)},\ \leavevmode\nobreak\ i=1,2$. Moreover their second moments are bounded, i.e.
$\displaystyle\mathbb{E}[||M_{n+1}^{(i)}||^{2}\ |\mathcal{I}_{n}]\leq K(1+||x_{n}||^{2}+||y_{n}||^{2}),\ \leavevmode\nobreak\ i=1,2$
for a suitable constant $K$.
Condition A.3.
Step-sizes $\{a(n)\}$ and $\{b(n)\}$ are positive scalars satisfying Robbins–Monro criteria
$\displaystyle\sum_{n}a(n)=\sum_{n}b(n)=\infty,\ \leavevmode\nobreak\ \sum_{n}(a(n)^{2}+b(n)^{2})<\infty,\ \leavevmode\nobreak\ \frac{b(n)}{a(n)}\longrightarrow 0$
In the above, $\frac{b(n)}{a(n)}\longrightarrow 0$ implies that the sequence $b(n)\longrightarrow 0$ converges to 0 faster compared to the sequence $\{a(n)\}$ and so the iterate in Equation (17) operates on a slower timescale compared to the iterate in Equation (16). Next, consider an o.d.e. defined as
$\displaystyle\dot{x}(t)=h(x(t),y)$ (22)
where $y$ is being held fixed and treated as a constant parameter.
Condition A.4.
The o.d.e. in Equation (22) has a globally asymptotically stable equilibrium $\lambda(y)$, where $\lambda:\mathbb{R}^{k}\longrightarrow\mathbb{R}^{d}$ is a Lipschitz map.
Given a small $\epsilon$, this condition tells us that the solution $x(t)$ tracks $\lambda(y(t))$ for $t>0$. This gives rise to the following o.d.e.:
$\displaystyle\dot{y}(t)=h(\lambda(y(t)),y(t)),$ (23)
where now the equation solely depends on trajectory $y(t)$. Finally, equation (23) yields another condition:
Condition A.5.
The o.d.e. (23) has a globally asymptotically stable equilibrium $y^{*}$.
Given the last two conditions, we expect the trajectory $(x(t),y(t))$ to approximately converge to the point $(\lambda(y^{*}),y^{*})$. The final condition necessary is the following.
Condition A.6.
Iterates $x_{n}$ and $y_{n}$ are almost surely bounded, i.e.
$\displaystyle\sup_{n}(||x_{n}||+||y_{n}||)<\infty,\ a.s.$
Now that we have all conditions in place, we present the Lemma in [11], guarantying asymptotic tracking of $\{\lambda(y_{n})\}$ by trajectory $\{x_{n}\}$ in the almost surely sense.
Lemma A.2 (Presented in [11]).
Given conditions A.1, A.2, A.3, A.4, A.5, A.6, the iterates converge almost surely to the optimum:
$\displaystyle(x_{n},y_{n})\overset{a.s.}{\longrightarrow}\{(\lambda(y),y)\ :\ y\in\mathbb{R}^{k}\}$
Proof.
The proof is omitted here as it can be found in [11]. ∎
Next, we present the main theorem, found in [10] or [11].
Theorem A.2 (Convergence of 2 timescale o.d.e. system [10]).
Given conditions A.1, A.2, A.3, A.4, A.5, A.6
$\displaystyle(x_{n},y_{n})\overset{a.s.}{\longrightarrow}(\lambda(y^{*}),y^{*})$
Proof.
The proof is omitted here as it can be found in [11]. ∎
In the remainder of this section, we show that all of the necessary assumptions hold for the newly defined iterates in Equations (20) and (21).
Lemma A.3 (Lipschitz continuity of $g_{c},g_{a}$).
Given the analytical gradients $g_{c},g_{a}$ defined above, let $(\omega_{t},\omega_{t}^{\prime})$ and $(\theta_{t},\theta_{t}^{\prime})$ be two sets of critic and actor parameters respectfully. Then,
$\displaystyle|g_{c}(\theta_{t},\omega_{t})-g_{c}(\theta_{t}^{\prime},\omega_{t}^{\prime})|$ $\displaystyle\leq L^{(c)}(|\theta_{t}-\theta_{t}^{\prime}|+|\omega_{t}-\omega_{t}^{\prime}|)$
$\displaystyle|g_{a}(\theta_{t},\omega_{t})-g_{a}(\theta_{t}^{\prime},\omega_{t}^{\prime})|$ $\displaystyle\leq L^{(a)}(|\theta_{t}-\theta_{t}^{\prime}|+|\omega_{t}-\omega_{t}^{\prime}|)$
for suitable Lipschitz constants $L^{(c)},L^{(a)}$.
Proof.
First, we focus on the gradient of the critic $g_{c}$. To show that $g_{c}$ is Lipschitz, it is sufficient to show that the partial derivatives $\partial\omega_{t}g_{c}$ and $\partial\theta_{t}g_{c}$ are continues and bounded. So,
$\displaystyle\partial\omega_{t}g_{c}$ $\displaystyle=\partial\omega_{t}\Big{|}\Big{|}\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{|}\Big{|}_{\theta_{t}}$
$\displaystyle=\Big{|}\Big{|}\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\Big{(}\nabla_{\omega_{t}}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}+\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}^{2}\delta_{\omega_{t}}\Big{)}\Big{|}\Big{|}_{\theta_{t}}$
$\displaystyle=\Big{|}\Big{|}\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\Big{(}\nabla_{\omega_{t}}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{)}\Big{|}\Big{|}_{\theta_{t}}$
because $\nabla_{\omega_{t}}^{2}V_{\omega_{t}}=0$. Applying claim A.1, it follows that $V_{\omega_{t}}$ and $\nabla_{\omega_{t}}V_{\omega_{t}}$ are continuous over the finite state space $\mathcal{S}$ (finite– thus bounded) for any $\omega_{t}$. Using the Extreme Value Theorem, it follows that, for a given parameter $\omega_{t}$, the functions $V_{\omega_{t}}$ and $\nabla_{\omega_{t}}V_{\omega_{t}}$ attain their maximums inside the bounded domain of the function, that is, the state space. Thus there exist bounds $B_{\omega_{t}}$ and $B_{\omega_{t}}^{\prime}$ such that $|V_{\omega_{t}}|<B_{\omega_{t}}$ and $|\nabla_{\omega_{t}}V_{\omega_{t}}|<B_{\omega_{t}}^{\prime}$. Moreover, Assumption 4.2, provides the existence of uniform bounds, denoted $B_{\omega},B_{\omega}^{\prime}$, such that $|V_{\omega_{t}}|<B_{\omega}$ and $|\nabla_{\omega_{t}}V_{\omega_{t}}|<B_{\omega}^{\prime}$. Furthermore, as stated above the reward $R_{t+1}$ is bounded by a constant $B_{R}$. Therefore, the term
$\displaystyle|\nabla_{\omega_{t}}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}|<B_{\partial\omega g_{c}}$
is bounded by constant $B_{\partial\omega g_{c}}\in\mathbb{R}$. Thus, it follows that the sum
$\displaystyle\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\Big{(}\nabla_{\omega_{t}}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{)}\leq\frac{B_{\partial\omega g_{c}}}{1-\gamma\lambda},\ \leavevmode\nobreak\ |\gamma\lambda|<1$
is bounded as well. From here we have that $\partial\omega_{t}g_{c}$ is bounded. The continuity of the partial derivative of the critic gradient follows from Claim A.1.
Similarly, taking the partial derivative $\partial\theta_{t}g_{c}$ gives:
$\displaystyle\partial\theta_{t}g_{c}$ $\displaystyle=\partial\theta_{t}\Big{|}\Big{|}\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{|}\Big{|}_{\theta_{t}}$
$\displaystyle=\partial\theta_{t}\sum_{s\in\mathcal{S}}\mu_{\theta_{t}}(s)\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}$
$\displaystyle=\sum_{s\in\mathcal{S}}\partial\theta_{t}\mu_{\theta_{t}}(s)\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}$
$\displaystyle\qquad\qquad\qquad+\sum_{s\in\mathcal{S}}\mu_{\theta_{t}}(s)\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\Big{(}\partial\theta_{t}\delta_{\omega_{t}}\nabla_{\omega_{t}}\delta_{\omega_{t}}+\delta_{\omega_{t}}\partial\theta_{t}\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{)}.$
First we focus on the $\sum_{s\in\mathcal{S}}\partial\theta_{t}\mu_{\theta_{t}}(s)$ term. The steady state distribution $\mu_{\theta_{t}}(s)$ depends on the parameter $\theta_{t}$ through the actor’s policy $\pi_{\theta_{t}}(\cdot|\cdot)$. By Assumption 4.1, the derivative of the policy is bounded. Hence $\partial\theta_{t}Pr(s\longrightarrow s^{\prime},k,\pi_{\theta_{t}})$ is bounded for all $k$ and $s\in\mathcal{S}$. It follows that the geometric series $\sum_{k=0}^{\infty}\gamma^{k}\partial\theta_{t}Pr(s\longrightarrow s^{\prime},k,\pi_{\theta_{t}})$ converges and therefore $\sum_{s\in\mathcal{S}}\partial\theta_{t}\mu_{\theta_{t}}(s)<\infty$. We know that the term $\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}$ is bounded as well. Finally, consider the terms $\partial\theta_{t}\delta_{\omega_{t}}$ and $\partial\theta_{t}\nabla_{\omega_{t}}\delta_{\omega_{t}}$. According to Assumption 4.1 and Claim A.1, the derivative of the feature vectors is bounded. Therefore, it follows that both $\partial\theta_{t}\delta_{\omega_{t}}$ and $\partial\theta_{t}\nabla_{\omega_{t}}\delta_{\omega_{t}}$ are bounded terms and hence $\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\Big{(}\partial\theta_{t}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}+\delta_{\omega_{t}}(t-K+l+1)\partial\theta_{t}\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{)}<\infty$. Therefore, we conclude that there exists a bound $B_{\partial\theta g_{c}}$ such that
$\displaystyle|\partial\theta_{t}g_{c}|<B_{\partial\theta g_{c}},\ \text{ for all }t.$
The continuity of the derivative of the gradient follows from the fact that all terms in the gradient are continuous. Next we show that given that the partial derivatives of the gradient are continuous and bounded, we have that the gradient itself is Lipschitz continues. Define
$\displaystyle S_{1}$ $\displaystyle=\sup_{\omega_{t}}\{|\partial\omega_{t}g_{c}(\theta_{t},\omega_{t})|\}$
$\displaystyle S_{2}$ $\displaystyle=\sup_{\theta_{t}}\{|\partial\theta_{t}g_{c}(\theta_{t},\omega_{t})|\}.$
Pick two sets of parameters $\theta_{t},\omega_{t}$ and $\theta_{t}^{\prime},\omega_{t}^{\prime}$. Then
$\displaystyle g_{c}(\theta_{t},\omega_{t})-g_{c}(\theta_{t}^{\prime},\omega_{t}^{\prime})$ $\displaystyle=g_{c}(\theta_{t},\omega_{t})-g_{c}(\theta_{t},\omega_{t}^{\prime})+g_{c}(\theta_{t},\omega_{t}^{\prime})-g_{c}(\theta_{t}^{\prime},\omega_{t}^{\prime})$
The triangle inequality implies:
$\displaystyle|g_{c}(\theta_{t},\omega_{t})-g_{c}(\theta_{t}^{\prime},\omega_{t}^{\prime})|\leq|g_{c}(\theta_{t},\omega_{t})-g_{c}(\theta_{t},\omega_{t}^{\prime})|+|g_{c}(\theta_{t},\omega_{t}^{\prime})-g_{c}(\theta_{t}^{\prime},\omega_{t}^{\prime})|.$
Then, since $g_{c}$ is continuous and the parameter spaces is bounded, by Assumptions 4.1, and 4.2, applying the Mean Value Theorem provides the existence of constants $c_{1}$ and $c_{2}$ such that
$\displaystyle\frac{g_{c}(\theta_{t},\omega_{t})-g_{c}(\theta_{t},\omega_{t}^{\prime})}{\omega_{t}-\omega_{t}^{\prime}}$ $\displaystyle=\partial\omega_{t}g_{c}(\theta_{t},c_{1})\leq S_{1}$
$\displaystyle\frac{g_{c}(\theta_{t},\omega_{t}^{\prime})-g_{c}(\theta_{t}^{\prime},\omega_{t}^{\prime})}{\theta_{t}-\theta_{t}^{\prime}}$ $\displaystyle=\partial\theta_{t}g_{c}(c_{2},\omega_{t})\leq S_{2}.$
Rearranging the terms in the two expressions above and substituting them back into the triangle inequality provides:
$\displaystyle|g_{c}(\theta_{t},\omega_{t})-g_{c}(\theta_{t}^{\prime},\omega_{t}^{\prime})|$ $\displaystyle\leq|g_{c}(\theta_{t},\omega_{t})-g_{c}(\theta_{t},\omega_{t}^{\prime})|+|g_{c}(\theta_{t},\omega_{t}^{\prime})-g_{c}(\theta_{t}^{\prime},\omega_{t}^{\prime})|$
$\displaystyle\leq S_{1}|\omega_{t}-\omega_{t}^{\prime}|+S_{2}|\theta_{t}-\theta_{t}^{\prime}|.$
Let $L^{(c)}=\max\{S_{1},S_{2}\}$. It follows that
$\displaystyle|g_{c}(\theta_{t},\omega_{t})-g_{c}(\theta_{t}^{\prime},\omega_{t}^{\prime})|$ $\displaystyle\leq L^{(c)}(|\omega_{t}-\omega_{t}^{\prime}|+|\theta_{t}-\theta_{t}^{\prime}|).$
Next we repeat the entire process from the derivations above to show that the Lipschitz condition holds for the actor gradient as well. Take the partial derivative with respect to critic parameter of the actor gradient $g_{a}$ to get
$\displaystyle\partial\omega_{t}g_{a}$ $\displaystyle=\partial\omega_{t}\mathbb{E}_{\pi_{\theta_{t}}}\Big{[}\nabla_{\theta_{t}}\ln\pi_{\theta_{t}}(A_{t}|S_{t})\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{]}$
$\displaystyle=\mathbb{E}_{\pi_{\theta_{t}}}\Big{[}\nabla_{\theta_{t}}\ln\pi_{\theta_{t}}(A_{t}|S_{t})\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\Big{(}\nabla_{\omega_{t}}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}+\delta_{\omega_{t}}(t+l)\nabla_{\omega_{t}}^{2}\delta_{\omega_{t}}\Big{)}\Big{]}$
$\displaystyle=\mathbb{E}_{\pi_{\theta_{t}}}\Big{[}\nabla_{\theta_{t}}\ln\pi_{\theta_{t}}(A_{t}|S_{t})\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\Big{(}\nabla_{\omega_{t}}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{)}\Big{]}.$
As before, all the terms inside the expectation are continuous and bounded and therefore there exists a constant $B_{\partial\omega g_{a}}$ such that
$\displaystyle|\partial\omega_{t}g_{a}(\theta_{t},\omega_{t})|\leq B_{\partial\omega g_{a}},\ \text{ for all }t.$
Furthermore, taking the partial derivative with respect to the actor parameter $\theta_{t}$, we get
$\displaystyle\partial\theta_{t}g_{a}$ $\displaystyle=\partial\theta_{t}\mathbb{E}_{\pi_{\theta_{t}}}\Big{[}\nabla_{\theta_{t}}\ln\pi_{\theta_{t}}(A_{t}|S_{t})\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{]}$
$\displaystyle=\partial\theta_{t}\sum_{s}\mu_{\theta_{t}}(s)\nabla_{\theta_{t}}\ln\pi_{\theta_{t}}(A_{t}|s)\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}$
$\displaystyle=\sum_{s}\partial\theta_{t}\mu_{\theta_{t}}(s)\nabla_{\theta_{t}}\ln\pi_{\theta_{t}}(A_{t}|s)\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}$
$\displaystyle+\sum_{s}\mu_{\theta_{t}}(s)\nabla^{2}_{\theta_{t}}\ln\pi_{\theta_{t}}(A_{t}|s)\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}$
$\displaystyle+\sum_{s}\mu_{\theta_{t}}(s)\nabla_{\theta_{t}}\ln\pi_{\theta_{t}}(A_{t}|s)\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\Big{(}\partial\theta_{t}\delta_{\omega_{t}}\nabla_{\omega_{t}}\delta_{\omega_{t}}+\delta_{\omega_{t}}\partial\theta_{t}\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{)}\Big{]}.$
Similarly to the critic’s first term, the first term is bounded and continuous. The second term is bounded and continuous due to Assumptions 4.1 and Claim 4.1 which tells us that $\nabla_{\theta_{t}}^{2}\ln\pi_{\theta_{t}}(A_{t}|S_{t})$ is bounded and continuous. Finally, the third term is bounded and continuous due to Claim A.1. Thus, there exists a constant $B_{\partial\theta g_{a}}$ such that
$\displaystyle|\partial\theta_{t}g_{a}(\theta_{t},\omega_{t})|\leq B_{\partial\theta g_{a}}.$
The rest of the derivation for finding the Lipshitz constant $L^{(a)}$ is along the similar lines as above. ∎
Therefore, we showed that Condition A.1 is satisfied. We proceed by working on Condition A.6.
Lemma A.4 (Bounded parameters).
The parameters $\omega_{t}$ and $\theta_{t}$ are uniformly bounded almost surely,
$\displaystyle\sup_{t}|\omega_{t}|+\sup_{t}|\theta_{t}|<\infty,\ a.s.$
Proof.
We show the derivation for the critic parameter. The derivation for the actor parameter follows the analogous argument. Consider the update scheme:
$\displaystyle\omega_{t+K}-\omega_{t+K-1}$ $\displaystyle=\beta(t+K-1)\Big{[}\frac{1}{K}\sum_{k=0}^{K-1}\Big{[}\sum_{l=k}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t+K-1}}(t+l)\nabla_{\omega_{t+K-1}}\delta_{\omega_{t+K-1}}(t+l)\Big{]}\Big{]}.$
Note that both $\delta_{\omega_{t+K-1}}(\cdot)$ and $\nabla_{\omega_{t+K-1}}\delta_{\omega_{t+K-1}}(\cdot)$ are bounded because the reward $R_{t+1}$ is bounded, and Claim A.1 and Assumption 4.2 both hold. Thus the following bound holds:
$\displaystyle|\omega_{t+K}-\omega_{t+K-1}|$ $\displaystyle\leq\beta(t+K-1)B_{\hat{g_{c}}}.$
Taking the limit of both sides, as $t\longrightarrow\infty$, we get
$\displaystyle\lim_{t\longrightarrow\infty}|\omega_{t+K}-\omega_{t+K-1}|$ $\displaystyle\leq\lim_{t\longrightarrow\infty}\beta(t+K-1)B_{\hat{g_{c}}}=0.$
Hence the sequence $\{\omega_{t}\}$ is Cauchy in the complete metric space and thus convergent. All convergent sequences are bounded and so we conclude that $\{\omega_{t}\}$ is bounded and hence almost surely bounded.
An analogous derivation follows for the sequence $\{\theta_{t}\}$. ∎
Next, we show that the terms $M_{t+1}^{(1)}$ and $M_{t+1}^{(2)}$ are martingale differences with bounded second moments.
Lemma A.5 (Martingale differences).
Objects $M_{t+1}^{(c)}$ and $M_{t+1}^{(a)}$ satisfy the martingale difference property and they have bounded second moments.
Proof.
Let $\mathcal{I}_{t}=\sigma(\omega_{t^{\prime}},\theta_{t^{\prime}},M_{t^{\prime}}^{(c)},M_{t^{\prime}}^{(a)},t^{\prime}\leq t)$ be a sigma algebra generated by $\omega_{t^{\prime}},\theta_{t^{\prime}},M_{t^{\prime}}^{(c)},M_{t^{\prime}}^{(a)}$, such that $t^{\prime}\leq t$. For the critic, we have that
$\displaystyle\mathbb{E}[M_{t+1}^{(c)}|\ \mathcal{I}_{t}]$ $\displaystyle=\mathbb{E}[\hat{g_{c}}(\theta_{t},\omega_{t})-g_{c}(\theta_{t},\omega_{t})|\ \mathcal{I}_{t}]=\mathbb{E}[\hat{g_{c}}(\theta_{t},\omega_{t})|\ \mathcal{I}_{t}]-\mathbb{E}[g_{c}(\theta_{t},\omega_{t})|\mathcal{I}_{t}]$
$\displaystyle=\mathbb{E}[\hat{g_{c}}(\theta_{t},\omega_{t})|\ \mathcal{I}_{t}]-\mathbb{E}\Big{[}\mathbb{E}_{s\sim\mu_{\theta_{t}}}\Big{[}\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{]}\Big{|}\mathcal{I}_{t}\Big{]}$
$\displaystyle=\mathbb{E}[\hat{g_{c}}(\theta_{t},\omega_{t})|\ \mathcal{I}_{t}]-\mathbb{E}\Big{[}\sum_{s}\mu_{\theta_{t}}(s)\Big{(}\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{)}\Big{|}\mathcal{I}_{t}\Big{]}.$
Consider again the filtration $\mathcal{I}_{t}=\sigma(\omega_{t^{\prime}},\theta_{t^{\prime}},M_{t^{\prime}}^{(c)},M_{t^{\prime}}^{(a)},t^{\prime}\leq t)$. Note that parameters $\omega_{t^{\prime}},\theta_{t^{\prime}}$ have been computed using a sample trajectory of states and actions. In other words, for parameters at time $t^{\prime}$, the agent observed a sample trajectory $\tau_{t^{\prime}-K:t^{\prime}}=(S_{t^{\prime}-K},A_{t^{\prime}-K},\dots,S_{t^{\prime}-1},A_{t^{\prime}-1},S_{t^{\prime}})$ in order to compute the gradients necessary for updates (5) and (7). Therefore, for parameter $\omega_{t}$, the agent observed a sample trajectory $\tau_{t-K:t}=(S_{t-K},A_{t-K},\dots,S_{t-1},A_{t-1},S_{t})$ and the finite sum inside of the inner expectation can be pulled out of the expectation as it is not a random variable with respect to filtration $\mathcal{I}_{t}$. To see it more clearly, note that $\delta_{\omega_{t}}(t-K+l+1)=R_{t-K+l+1}+\gamma V_{\omega_{t}}(S_{t-K+l+1})-V_{\omega_{t}}(S_{t-K+l})$ for $l=0,\dots,K-1$. So, when $l=K-1$, we get $\delta_{\omega_{t}}(t)=R_{t}(S_{t-1},A_{t-1})+\gamma V_{\omega_{t}}(S_{t})-V_{\omega_{t}}(S_{t-1})$ which has been observed along trajectory $\tau_{t-K:t}=(S_{t-K},A_{t-K},\dots,S_{t-1},A_{t-1},S_{t})$. Thus,
$\displaystyle\mathbb{E}[M_{t+1}^{(c)}|\ \mathcal{I}_{t}]$ $\displaystyle=\mathbb{E}[\hat{g_{c}}(\theta_{t},\omega_{t})|\ \mathcal{I}_{t}]-\mathbb{E}\Big{[}\sum_{s}\mu_{\theta_{t}}(s)\Big{(}\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{)}\Big{|}\mathcal{I}_{t}\Big{]}$
$\displaystyle=\mathbb{E}[\hat{g_{c}}(\theta_{t},\omega_{t})|\ \mathcal{I}_{t}]-\mathbb{E}\Big{[}\sum_{l=0}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t}}(t-K+l+1)\nabla_{\omega_{t}}\delta_{\omega_{t}}\Big{|}\mathcal{I}_{t}\Big{]}=0.$
The analogous argument is made in the case of actor iterates.
Next, we want to show that the second moment is bounded for the critic iterate. Given the derived bounds in Lemma A.3, it follows that
$\displaystyle\mathbb{E}[|M_{t+1}^{(c)}|^{2}|\ \mathcal{I}_{t}]$ $\displaystyle=\mathbb{E}[|\hat{g_{c}}(\theta_{t},\omega_{t})-g_{c}(\theta_{t},\omega_{t})|^{2}|\ \mathcal{I}_{t}]\leq 2\mathbb{E}[|\hat{g_{c}}(\theta_{t},\omega_{t})|^{2}+|g_{c}(\theta_{t},\omega_{t})|^{2}|\ \mathcal{I}_{t}]$
$\displaystyle=2\mathbb{E}[|\hat{g_{c}}(\theta_{t},\omega_{t})|^{2}|\ \mathcal{I}_{t}]+2\mathbb{E}[|g_{c}(\theta_{t},\omega_{t})|^{2}|\ \mathcal{I}_{t}]$
$\displaystyle\leq 2\mathbb{E}[|\hat{g_{c}}(\theta_{t},\omega_{t})|^{2}|\ \mathcal{I}_{t}]+2L^{(c)^{2}}\mathbb{E}[(|\theta_{t}|+|\omega_{t}|)^{2}|\mathcal{I}_{t}]$
$\displaystyle\leq 2\mathbb{E}[|\hat{g_{c}}(\theta_{t},\omega_{t})|^{2}|\ \mathcal{I}_{t}]+4L^{(c)^{2}}\mathbb{E}[|\theta_{t}|^{2}+|\omega_{t}|^{2}|\mathcal{I}_{t}]$
$\displaystyle\leq 2B_{\hat{g_{c}}}+4L^{(c)^{2}}(|\theta_{t}|^{2}+|\omega_{t}|^{2})$
where, as in the lemma A.4, the approximate gradient $\hat{g_{c}}(\theta_{t},\omega_{t})$ is bounded with bound $B_{\hat{g_{c}}}$. Let $B_{M^{(c)}}:=\max\{2B_{\hat{g_{c}}},4L^{(c)^{2}}\}$. It follows that
$\displaystyle\mathbb{E}[|M_{t+1}^{(c)}|^{2}|\ \mathcal{I}_{t}]\leq B_{M^{(c)}}(1+|\theta_{t}|^{2}+|\omega_{t}|^{2}).$
An analogous derivation holds for $M_{t+1}^{(a)}$. ∎
Therefore, condition A.2 is satisfied as well.
Lemma A.6 (Step size choice).
Denote two real valued, decreasing sequences by $\alpha(t),\beta(t)$ respectfully. Then, we define them as
$\displaystyle\alpha(t)$ $\displaystyle=\frac{1}{t^{2/3}},\ t=1,2,\dots$
$\displaystyle\beta(t)$ $\displaystyle=\frac{1}{t},\ t=1,2,\dots.$
It follows that $\alpha(t)$ and $\beta(t)$ satisfy the Robbins-Monro criterion.
Proof.
Note that for discrete step sizes $t$, it follows that
$\displaystyle\sum_{t}\alpha(t)=\sum_{t}\beta(t)=\infty$
$\displaystyle\sum_{t}\alpha(t)^{2}<\infty,\ \sum_{t}\beta(t)^{2}<\infty\ \sum_{t}(\alpha(t)^{2}+\beta(t)^{2})<\infty$
$\displaystyle\frac{\beta(t)}{\alpha(t)}=\frac{1}{t^{1/3}}\longrightarrow 0.$
Therefore, Condition A.3 is satisfied as well. Next consider the mean ODE, defined as
$\displaystyle\dot{\omega}(t)=g_{c}(\theta,\omega(t))$ (24)
for a fixed $\theta$. We want to show that Equation (24) has a unique asymptotically stable equilibrium point $\lambda(\theta)$ where $\lambda$ is a Lipschitz map.
Lemma A.7 (Asymptotic stability of $\dot{\omega}(t)$).
Fix the iterate $\theta_{t}\equiv\theta_{0}$ and consider the mean ordinary differential equation $\dot{\omega_{t}}$, defined as
$\displaystyle\dot{\omega}(t)=g_{c}(\theta_{0},\omega(t)).$ (25)
It follows that there exists a unique asymptotically stable equilibrium point $\omega(\theta_{0})^{*}$.
Proof.
Because $g_{c}$ is Lipschitz, it follows that the ODE (24) is well posed, i.e. it has a unique solution for each initial condition $(\theta_{0},\omega(0))$ and the solution continuously depends on $(\theta_{0},\omega(0))$. By the Picard-Lindelöf theorem, a unique equilibrium point of the system exists on some interval $I=[t_{0}-\epsilon,t_{0}+\epsilon]$. Denote such an equilibrium point as $\omega(\theta_{0})^{*}$, and define a corresponding Lypaunov function $\mathcal{L}_{\omega}(\omega(t)-\omega(\theta_{0})^{*})$ as
$\displaystyle\mathcal{L}_{\omega}(\omega(t)-\omega(\theta_{0})^{*}):=L_{g_{c}}(\theta,\omega(\theta_{0})^{*})-L_{g_{c}}(\theta,\omega(t)).$
Computing the derivative of $\mathcal{L}_{\omega}$ with respect to $t$ provides:
$\displaystyle\frac{d}{dt}\mathcal{L}_{\omega}(\omega(t)-\omega(\theta_{0})^{*})$ $\displaystyle=\frac{d}{d\omega}\mathcal{L}_{\omega}(\omega(t)-\omega(\theta_{0})^{*})\frac{d}{dt}\omega(t)=-\frac{\partial}{\partial\omega}L_{g_{c}}(\theta,\omega(t))\cdot g_{c}(\theta,\omega(t))=-g_{c}(\theta,\omega(t))^{2}.$
Because $L_{g_{c}}$ is convex, it follows that the Lyapunov function $\mathcal{L}_{\omega}$ is positive definite. Furthermore, it is continuous differentiable according to Claim A.1. Finally, it satisfies
$\displaystyle\frac{d}{d\omega}\mathcal{L}_{\omega}(\omega(t)-\omega(\theta_{0})^{*})\frac{d}{dt}\omega(t)\leq 0,\ \text{ for all }\omega(t).$
Then, by the Lyapunov theorem on stability, it follows that $\omega(\theta_{0})^{*}$ is a Lyapunov stable equilibrium of our mean ODE. Furthermore, because ODE (24) is well posed for any initial solution and it is Lyapunov stable, it follows that ODE (24) is (globally) asymptotically stable with the stable equilibrium $\omega(\theta_{0})^{*}$. ∎
We also need to establish the existence of a Lipschitz map $\lambda$ which maps from the space of parameters $\theta$ to the space of parameters $\omega$. In order to this, we follow the work of [24].
Lemma A.8 (Existence of Lipschitz map).
There exists a map $\lambda(\theta):\Theta\longrightarrow\Omega$ such that $\lambda(\cdot)$ is Lipschitz.
Proof.
We proceed by applying the implicit function theorem (IFT) to the critic’s gradient and establish that there exists a continuously differentiable function $\lambda(\cdot)$. Consider function $f$ defined as
$\displaystyle f(\theta,\cdot):=\nabla_{\omega}L_{g_{c}}(\theta,\cdot).$
Let $\omega(\theta_{0})^{*}$ be the critical point of function $f$ such that
$\displaystyle f(\theta_{0},\omega(\theta_{0})^{*})=\nabla_{\omega}L_{g_{c}}(\theta_{0},\omega(\theta_{0})^{*})=0.$
Let $(\theta_{0},\omega_{0})$ be the initial parameter pair and define a neighbourhood $U_{\theta_{0}}(\omega_{0})$ that connects $\omega_{0}$ with the local minimum $\omega(\theta_{0})^{*}$ of function $f$ such that there are no other local minima in this neighbourhood. This can be done by defining a neighbourhood as a narrow strip on the loss surface of function $L_{g_{c}}$ which starts around $\omega_{0}$ and extends towards the local minima $\omega(\theta_{0})^{*}$. Then by the implicit function theorem (IFT), it follows that there exists a neighbourhood $V(\theta_{0})$ around initial point $\theta_{0}$ and continuously differentiable function $\lambda:V\longrightarrow U_{\theta}$ such that
$\displaystyle f(\theta_{0},\lambda(\theta))=\nabla_{\omega}L_{g_{c}}(\theta_{0},\omega(\theta_{0})^{*})=0,\ \theta\in V(\theta_{0}).$
where
$\displaystyle\lambda(\theta)=\omega(\theta_{0})^{*},\ \theta\in V(\theta_{0}).$
Because the parameter spaces $\Theta$ and $\Omega$ are bounded, it follows that by Extreme Value theorem, that function $\lambda$ is bounded. Furthermore, because $\lambda$ is continues, differentiable and bounded, it follows that by the Mean Value theorem that $\lambda$ is Lipschitz. ∎
Now that we showed that condition A.1 is satisfied, it follows that by applying lemma A.2, we establish convergence of the system
$\displaystyle\dot{\omega}(t)$ $\displaystyle=g_{c}(\omega(t),\theta(t))$
$\displaystyle\dot{\theta}(t)$ $\displaystyle=0$
which implies that
$\displaystyle(\omega_{t},\theta_{t})\overset{a.s.}{\longrightarrow}\{(\lambda(\theta),\theta)\ :\ \theta\in\mathbb{R}^{n}\}.$ (26)
The last condition we need to check is Condition A.5. In this case, we cannot apply the Lyapunov technique again as the policy gradient not need be a positive definite function. However, we consider the following well established result about gradient systems.
Lemma A.9 (Convergence of gradient trajectory [1]).
Assume that $f$ is a $C^{2}$ function and let $x(t)$ be a solution trajectory of the gradient system
$\displaystyle\dot{x}(t)=-\nabla f(x(t))$
contained in a compact set $K\subseteq\mathbb{R}^{n}$. Then $x(t)$ approaches critical set $C_{K}=\{y\in K\ :\ \nabla f(y)=0\}$ as $t$ approaches infinity. That is, $\lim_{t\longrightarrow\infty}||x(t)-y||=0$.
Proof.
The proof is shown by using Theorem 9.4.4 in [23] and Lemma 3.1 in [30]. ∎
Lemma A.10 (Asymptotic stability of $\dot{\theta}(t)$).
Consider the mean ordinary differential equation $\dot{\theta}(t)$, defined as
$\displaystyle\dot{\theta}(t)=g_{a}\big{(}\theta(t),\lambda(\theta(t))\big{)}.$ (27)
It follows that there exists a unique asymptotically stable equilibrium point $\big{(}\theta^{*},\lambda(\theta^{*})\big{)}$.
Proof.
In the case of the ODE $\dot{\theta}(t)$ above, the dynamics from Lemma A.9 is the negative of the actor’s loss function, defined as $-f(\theta(t)):=L_{a}(\theta(t))$. Furthermore, we approximation the gradient with the projected gradient $g_{a}(\theta(t),\lambda(\theta(t)))$, as the projection error is orthogonal to the vector space $\Psi_{\theta}$, and get
$\displaystyle g_{a}(\theta(t),\lambda(\theta(t)))\approx\nabla L_{a}(\theta(t))=-\nabla f(\theta(t)).$
Hence, we consider the system
$\displaystyle\dot{\theta}(t)=g_{a}\big{(}\theta(t),\lambda(\theta(t))\big{)}.$
According to claim 4.1, $L_{a}$ is twice continuously differentiable with respect to actor parameter $\theta$. Furthermore, the trajectory $\big{(}\theta(t),\lambda(\theta(t))\big{)}$ remains in the compact space $K$ due to the parameter space $\Omega\times\Theta$ being closed and bounded. Hence, by applying lemma A.9, it follows that
$\displaystyle\big{(}\theta(t),\lambda(\theta(t))\big{)}\overset{a.s.}{\longrightarrow}\big{(}\theta^{*},\lambda(\theta^{*})\big{)}.$
Finally, due to conditions A.1, A.2, A.3, A.4, A.5, A.6 being satisfied, we can apply Lemma A.2 and Theorem A.2 to show that the Actor-Critic algorithm converges to its approximate optimal policy $\pi^{*}$.
This completes the proof of Theorem 5.1.
A.9 Further details on the experiments
The sampling procedure is done using a standard CPU offered by Google Colab, although we note that some speed up might be possible by leveraging GPU functionality. For the 100100 100100-node graph, with edge probability of $p=0.3$, which corresponds to fiber point $g\in\mathbb{Z}^{4950}$, the Actor-Critic move generator Algorithm 2 approximately takes 4.54.5 4.54.5 minutes to generate 150 feasible moves.
Data generation.
We test the algorithm on a co-authorship dataset and on synthetic data. The co-authorship data set was published in [27]. Since this is a sparse network of a very particular structure (see Figure 2), we also test our algorithm on families of randomly generated graphs. Here we show results obtained by generating Erdös-Rényi graphs with edge probability parameter $p$. We use https://networkx.org package in python for this step. We choose these types of networks because they are easy to generate and allow sparsity control in the data, in order to gain a better understanding of algorithm robustness.
A.10 Pseudo code for Actor-Critic Move Generator
Given an initial point $G$ (e.g. $G$ is a graph), Algorithm 1 decomposes $G$ into sub-points for which we can compute the lattice basis. In the case of graph data, this decomposition can be done using $k$-core decomposition, bridge cuts or induced subgraphs.
Algorithm 1 Initial Point Decomposition
Input: Initial point (e.g. Graph $G=(E,V)$)
      Output: Collection of sub-initial solutions $\mathcal{G}$, $\mathcal{B}_{\mathcal{G}}$ lattice basis for each sub-problem in $\mathcal{G}$
1:Use k-core decomposition, bridge cuts, induced subgraphs to obtain $\mathcal{G}$
2:for $g\in\mathcal{G}$ do
3:     Compute sufficient statistic $Ag=b_{g}$, lattice basis $\mathcal{B}_{g}$
4:end for
After obtaining sub-problems and their corresponding lattice basis, we employ the A2C algorithm to generate moves on the sub-problems.
Algorithm 2 Pseudo code of A2C move generator
Input: Learning rates $\alpha,\beta$, Collection of subgraphs $\mathcal{G}$, Lattice basis $\mathcal{B}_{\mathcal{G}}$, Size of roll-out $K$, Total number of steps $t_{max}$
      Output: $\pi^{*}$ OR/AND set of feasible moves $\mathcal{M}$
1:repeat
2:     Reset gradients: $d\theta\leftarrow 0\ ,d\omega\leftarrow 0$
3:     $t_{start}\leftarrow t$
4:     Get state $S_{t}=s$
5:     repeat
6:         Preform $A_{t}$ according to policy $\pi(A_{t}|S_{t};\theta)$
7:         Receive reward $R_{t+1}$ and new state $S_{t+1}$
8:         $t\leftarrow t+1$
9:     until terminal $S_{t}$ or $t-t_{start}==K$
10:     Compute temporal differences for each time step during roll-out
$\displaystyle\delta_{\omega_{t+K-1}}(t)=R_{t+1}(S_{t},A_{t})+\gamma V_{\omega_{t+K-1}}(S_{t}+A_{t})-V_{\omega_{t+K-1}}(S_{t}),\ t=t_{start},\dots,K$
11:     Accumulate gradients w.r.t. $\theta$ and update parameter:
$\displaystyle\theta_{t+K}$ $\displaystyle=\theta_{t+K-1}+\alpha(t+K-1)\Big{[}\frac{1}{K}\sum_{k=0}^{K-1}\nabla_{\theta_{t+K-1}}\ln\pi_{\theta_{t+K-1}}(A_{t+k}|S_{t+k})\sum_{l=k}^{K-1}(\gamma\lambda)^{l-k}\delta_{\omega_{t+K-1}}(t+l)\Big{]}$
12:     Accumulate gradients w.r.t. $\omega$ and update parameter:
$\displaystyle\omega_{t+K}$ $\displaystyle=\omega_{t+K-1}+\beta(t+K-1)\Big{[}\frac{1}{K}\sum_{k=0}^{K-1}\Big{[}\sum_{l=k}^{K-1}(\gamma\lambda)^{l}\delta_{\omega_{t+K-1}}(t+l)\nabla_{\omega_{t+K-1}}\delta_{\omega_{t+K-1}}(t+l)\Big{]}\Big{]}$
13:until $t>t_{max}$
Next we apply Algorithm 3 to the generated moves in order to lift them into a higher dimensional space. This space corresponds to the size of the original problem we set out to sample. Each sub-problem and its generated moves hold information about which sub-structure they belong to. We track the edge labels of the sub-moves and compare them to the edge labels of the original problem. Our lifted move starts as an empty list which iteratively gets populated with integer values. Whenever we have a match between edges from the the sub-move and original problem, we append the value of the sub-problem at that position to the empty list. If there is no such match, we append a zero. In the end, we obtain a move with size matching the original problem. Because we are padding the sub-move with zeros, such move is still applicable to the original problem.
Algorithm 3 Move Permutation
Input: Feasible moves $\mathcal{M}$, list of lexicographically ordered edges $\mathcal{E}$ of fully connected graph
      Output: Permuted moves $\mathcal{M}^{*}$
1:for $m\in\mathcal{M}$ do
2:     $m^{*}\leftarrow[\ ]$
3:     for $e\text{ in }\mathcal{E}$ do
4:         if $e\text{ not in }m$ then
5:              $m^{*}.\text{append}(0)$
6:         else
7:              $m^{*}.\text{append}(m[indx(e)])$
8:         end if
9:     end for
10:end for
Finally, having obtained the permuted moves $\mathcal{M}^{*}$ from Algorithm 3, we proceed to compute the exchangeable sample $\mathcal{D}_{A}$ and all possible $\mathcal{S}^{0}$ states that can be reached by applying them. The sampling proceeds by uniformly sampling a move from the set of all permuted moves and applying it to the current point in the fiber. If the next point in the fiber is feasible, this point becomes the current point. Moreover, we add the sampled point to $\mathcal{S}^{0}$ if it is already not there.
Algorithm 4 Random Sampling
Input: List of permuted moves $\mathcal{M}^{*}$, number of samples $N$, length of Markov Chain $L$, initial point $s_{0}$
      Output: Exchangeable sample $\mathcal{D}_{A}$ AND/OR all discovered states $\mathcal{S}^{0}$
1:$\mathcal{D}_{A}\leftarrow[\ ]$
2:$\mathcal{S}^{0}\leftarrow[\ ]$
3:for $n\in\textbf{range}(N)$ do
4:     $s\leftarrow s_{0}$
5:     for $l\in\textbf{range}(L)$ do
6:         $\text{indx}\leftarrow\text{uniform}[0,len(\mathcal{M}^{*})]$
7:         $m\leftarrow\mathcal{M}^{*}[\text{indx}]$
8:         $s^{\prime}\leftarrow s+m$
9:         if $s^{\prime}-$feasible then
10:              $s\leftarrow s^{\prime}$
11:              if $s\text{ not in }\mathcal{S}^{0}$ then
12:                  $\mathcal{S}^{0}.\text{append}(s)$
13:              end if
14:         end if
15:     end for
16:     $\mathcal{D}_{A}.\text{append}(s)$
17:end for